### v1

# Polynomial invariants for $3$-dimensional Leibniz algebras 111 The work is supported by the FCT 2023.08031.CEECIND, UIDB/00212/2020, UIDP/00212/2020, and FAPESP 2023/17918-2

Abstract: For every $3$-dimensional non-Lie Leibniz algebras over complex numbers we describe the algebra of polynomial invariants and the group of automorphisms. As a consequence, we establish that two non-nilpotent $3$-dimensional non-Lie Leibniz algebras can be distinguished by means of the traces of degrees $\leqslant 2$ and the dimensions of groups of automorphisms.
Keywords: Leibniz algebras, polynomial invariants, separating set, generating set.
MSC2020: 13A50, 15A72, 17A32, 17A36, 20F29.
1 Introduction
1.1 Algebra of invariants
Assume that $\mathbb{F}$ is an algebraically closed field of an arbitrary characteristic $\mathop{\rm char}{\mathbb{F}}$. All vector spaces and algebras are over $\mathbb{F}$.
Consider an algebra $\mathcal{A}$ is of dimension $n$, i.e., $\mathcal{A}$ is a vector space with a bilinear multiplication, and a subgroup $G$ of the group of all automorphisms $\mathop{\rm Aut}(\mathcal{A})\leqslant{\rm GL}_{n}$ of $\mathcal{A}$. The group $G$ naturally acts on $\mathcal{A}^{m}=\mathcal{A}\oplus\cdots\oplus\mathcal{A}$ ($m$ times) diagonally: $g{\underline{a}}=(g\cdot a_{1},\ldots,g\cdot a_{m})$ for each $g\in G$ and ${\underline{a}}=(a_{1},\ldots,a_{m})$ from $\mathcal{A}^{m}$. The coordinate ring of the vector space $\mathcal{A}^{m}$ is a polynomial (i.e. commutative and associative) algebra
[…]
Fix some basis $\{e_{1},\ldots,e_{n}\}$ for $\mathcal{A}$, and for every $a\in\mathcal{A}$ denote by $(a)_{i}$ the $i-{\rm th}$ coordinate of $a$ with respect the given basis. We can consider elements of $\mathbb{F}[\mathcal{A}^{m}]$ as polynomial functions $\mathcal{A}^{m}\to\mathbb{F}$, since $x_{ri}$ can be interpreted as a function $\mathcal{A}^{m}\to\mathbb{F}$ given by $x_{ri}({\underline{a}})=(a_{r})_{i}$. The action of $G$ on $\mathcal{A}^{m}$ induces the action of the coordinate ring $\mathbb{F}[\mathcal{A}^{m}]$ as follows: $(g\cdot f)({\underline{a}})=f(g^{-1}\cdot{\underline{a}})$ for all $g\in G$, $f\in\mathbb{F}[\mathcal{A}^{m}]$ and ${\underline{a}}\in\mathcal{A}^{m}$. The algebra of $G$-invariants of the $m$-tuple of the algebra $\mathcal{A}$ is
$\mathbb{F}[\mathcal{A}^{m}]^{G}=\{f\in\mathbb{F}[\mathcal{A}^{m}]\,|\,gf=f\text{ for all }g\in G\},$
or, equivalently,
$\mathbb{F}[\mathcal{A}^{m}]^{G}=\{f\in\mathbb{F}[\mathcal{A}^{m}]\,|\,f(g{\underline{a}})=f({\underline{a}})\text{ for all }g\in G,\;{\underline{a}}\in\mathcal{A}^{m}\}.$
For short, this algebra of $\mathop{\rm Aut}(\mathcal{A})$-invariants of the $m$-tuple of $\mathcal{A}$ is called the algebra of invariants of the $m$-tuple of $\mathcal{A}$, and we denote it by
$I_{m}(\mathcal{A}):=\mathbb{F}[\mathcal{A}^{m}]^{\mathop{\rm Aut}(\mathcal{A})}.$
It is well known that the so-called operator traces (see Section 2.2 for the details) are $\mathop{\rm Aut}(\mathcal{A})$-invariants. Denote by $\mathop{\rm Tr}(\mathcal{A})_{m}$ the subalgebra of $I_{m}(\mathcal{A})$ generated by all operator traces together with $1$. We say that the Artin–Procesi–Iltyakov Equality holds for $\mathcal{A}^{m}$ if
$I_{m}(\mathcal{A})=\mathop{\rm Tr}(\mathcal{A})_{m}.$
The Artin–Procesi–Iltyakov Equality is known to be valid for $\mathcal{A}^{m}$ with $m\geqslant 1$ for the following algebras:
  1. $\bullet$
    $\mathcal{A}=M_{n}$ is the algebra of all $n\times n$ matrices over $\mathbb{F}$, in case $\mathop{\rm char}{\mathbb{F}}=0$ or $\mathop{\rm char}{\mathbb{F}}>n$. Generators for the algebra of invariants $I_{m}(M_{n})$ were independently described in [36] and [32] in the case of charactristic zero, and in [11] in the case of positive characteristic. Minimal generating sets for $I_{m}(M_{2})$ and $I_{m}(M_{3})$ were obtained in [33, 10, 25, 26, 27] and for $I_{m}(M_{4})$ and $I_{m}(M_{5})$ with small values of $m$, in [39, 12, 9], assuming $\mathop{\rm char}\mathbb{F}=0$.
  2. $\bullet$
    $\mathcal{A}=\mathbf{O}$ is the octonion algebra, in case $\mathop{\rm char}{\mathbb{F}}\neq 2$. Generators for the algebra of invariants $I_{m}(\mathbf{O})$ were constructed in [35] over the field of complex numbers $\mathbb{C}$ and in [40] over an arbitrary infinite field of odd characteristic. In case $\mathop{\rm char}{\mathbb{F}}\neq 2$, a minimal generating set was constructed [28], using the classification of pairs of octonions from [29].
  3. $\bullet$
    $\mathcal{A}={\mathbb{A}}$ is the split Albert algebra, i.e., the exceptional simple Jordan algebra of $3\times 3$ Hermitian matrices over $\mathbf{O}$ with the symmetric multiplication $a\circ b=(ab+ba)/2$, in case $\mathop{\rm char}{\mathbb{F}}=0$ and $m\in\{1,2\}$ (see [16, 31]).
More details about the Artin–Procesi–Iltyakov Equality can be found in [1].
1.2 Leibniz algebras
An algebra $\mathcal{L}$ is called a (right) Leibniz algebra if it satisfies the identity
$(xy)z=(xz)y+x(yz).$
Leibniz algebras present a ”non antisymmetric” generalization of Lie algebras. It was at first introduced by A. Bloh [3] and then independently by J.-L. Loday [23]. Recently, they appeared in many geometric and physics applications (see, for example, [4, 22, 21, 37] and references therein). A systematic study of algebraic properties of Leibniz algebras was started by J.-L. Loday and T. Pirashvili [24]. So, several classical theorems from Lie algebras theory have been extended to the Leibniz algebras case; many classification results regarding nilpotent, solvable, etc. Leibniz algebras have been obtained (see, for example, [13, 2, 7, 6, 14, 20, 19, 30, 38] and references therein); for results on algebraic and geometric classifications of low dimensional Leibniz algebras see [17, 18] and references therein.
Given a Leibniz algebra $\mathcal{L}$, we denote by ${\rm Leib}(\mathcal{L})$ the $\mathbb{F}$-span of $a^{2}$ for all $a\in\mathcal{L}$. Since
$yx^{2}=0\;\text{ and }\;x^{2}y=(x^{2}+y)^{2}-(x^{2})^{2}-y^{2},$
we have that ${\rm Leib}(\mathcal{L})$ is an ideal of $\mathcal{L}$. Note that ${\rm Leib}(\mathcal{L})$ is non-zero if and only if $\mathcal{L}$ is not a Lie algebra. On the other hand, if ${\rm Leib}(\mathcal{L})=\mathcal{L}$, then $\mathcal{L}$ has zero multiplication. Therefore, non-Lie Leibniz algebra is never simple.
1.3 Results
In this section we assume that $\mathbb{F}=\mathbb{C}$ is the field of complex numbers. Denote by ${\bf L}_{3}$ the family of all $3$-dimensional non-Lie Leibniz algebras over the field $\mathbb{F}$. The classification of algebras from ${\bf L}_{3}$ modulo the action of the group of automorphisms was independently obtained by I.M. Rikhsiboev, I.S. Rakhimov [34] and by J.M. Casas, M.A. Insua, M. Ladra, S. Ladra [5]. The equivalence of these classifications was pointed out in the Remark at the end of paper [34]. In our paper, we use the classification from [34] (see Theorem 9).
In Section 2, we provide key definitions of generic elements and operator traces. In Section 3, we present an explicit formula for calculating some operator traces (see Proposition 4) and describe the reduction of the general case to multilinear in characteristic zero (see Proposition 6).
Assume $\mathcal{L}\in{\bf L}_{3}$. In Sections 46 we assume that $\mathbb{F}=\mathbb{C}$ is the field of complex numbers. In Section 4 we explicitly describe the group $\mathop{\rm Aut}(\mathcal{L})$ of all automorphisms of $\mathcal{L}$ (see Theorem 10). Then in Section 5 we obtain a generating set for the algebra of invariants $I_{m}(\mathcal{L})$ (see Theorem 14). As a consequence, we describe when Artin–Procesi–Iltyakov Equality holds for $\mathcal{L}$ (see Corollary 16). In Corollary 18 we obtain $I_{m}(\mathcal{L})=\mathbb{F}$ for all $m\geqslant 1$ if and only if $\mathcal{L}$ is nilpotent. As a consequence, in Corollary 19 we show that any two non-nilpotent algebras from ${\bf L}_{3}$ can be distinguished by means of
  1. $\bullet$
    the traces of degrees $\leqslant 2$,
  2. $\bullet$
    the dimensions of its groups of automorphisms.
Note that the problem of distinguishing low-dimensional algebras using their algebraic properties has recently been studied. Specifically, in [8], it was shown that any two-dimensional Jordan algebras over a finite field $\mathbb{F}$ with odd characteristic are isomorphic if and only if they satisfy the same polynomial identities. In [15], it was established that polynomial identities distinguish non-associative two-dimensional Novikov algebras over the complex numbers.
1.4 Notations
A monomial $w=x_{r_{1},i_{1}}\cdots x_{r_{k},i_{k}}$ from $\mathbb{F}[\mathcal{A}^{m}]$ has multidegree $\mathop{\rm mdeg}(w)=(\delta_{1},\ldots,\delta_{m})\in\mathbb{N}^{m}$, where $\delta_{r}$ is the number of letters of $w$ lying in the set $\{x_{r1},\ldots,x_{rn}\}$ and =$\mathbb{N}=\{0,1,2,\ldots\}$. As an example, $\mathop{\rm mdeg}(x_{31}x_{32}x_{21}x_{12})=(1,1,2)$ for $m=3$. For short, we denote the multidegree $(1,\ldots,1)\in\mathbb{N}^{m}$ by $1^{m}$. If $f\in\mathbb{F}[\mathcal{A}^{m}]$ is a linear combination of monomials of the same multidegree ${\underline{d}}$, then we say that $f$ is $\mathbb{N}^{m}$-homogeneous of multidegree ${\underline{d}}$. In other words, we have defined the $\mathbb{N}^{m}$-grading of $\mathbb{F}[\mathcal{A}^{m}]$ by multidegrees. Since $\mathbb{F}$ is infinite, the algebra of invariants $\mathbb{F}[\mathcal{A}^{m}]^{G}$ also has the $\mathbb{N}^{m}$-grading by multidegrees. An $\mathbb{N}$-homogeneous element $f\in\mathbb{F}[\mathcal{A}^{m}]$ of multidegree $(d_{1},\ldots,d_{m})$ with $d_{1},\ldots,d_{m}\in\{0,1\}$ is called multilinear. For short, we denote by $x_{r_{1},i_{1}}\cdots\widehat{x_{r_{l},i_{l}}}\cdots x_{r_{k},i_{k}}$ the monomial $x_{r_{1},i_{1}}\cdots x_{r_{l-1},i_{l-1}}x_{r_{l+1},i_{l+1}}\cdots x_{r_{k},i_{k}}$ from $\mathbb{F}[\mathcal{A}^{m}]$, where $1\leqslant l\leqslant k$. Given a subset $S\subset\mathcal{A}$, we denote by $\mathop{\rm alg}\{S\}$ the subalgebra of $\mathcal{A}$ (without unity in general) generated by $S$. We write $\mathbb{F}^{\times}$ for the set of non-zero elements of $\mathbb{F}$.
If there exists $k\geqslant 1$ such that for every $a_{1},\ldots,a_{k}\in\mathcal{A}$ each product of $a_{1},\ldots,a_{k}$ is zero in $\mathcal{A}$, then the algebra $\mathcal{A}$ is called nilpotent and the least possible $k\geqslant 1$ with the above property is called the nilpotency class ${\rm ncl}(\mathcal{A})$ for $\mathcal{A}$; otherwise, we write ${\rm ncl}(\mathcal{A})=\infty$.
2 Preliminaries on invariants
2.1 Generic elements
To explicitly define the action of $G$ on $\mathbb{F}[\mathcal{A}^{m}]$ consider the algebra $\widehat{\mathcal{A}}_{m}=\mathcal{A}\otimes_{\mathbb{F}}\mathbb{F}[\mathcal{A}^{m}]$, which is a $G$-module, where the multiplication and the $G$-action are defined as follows: $(a\otimes f)(b\otimes h)=ab\otimes fh$ and $g\bullet(a\otimes f)=g\cdot a\otimes f$ for all $g\in G$, $a,b\in\mathcal{A}$ and $f,h\in\mathbb{F}[\mathcal{A}^{m}]$. Define by
$X_{r}=\left(\begin{array}[]{c}x_{r1}\\ \vdots\\ x_{rn}\\ \end{array}\right):=e_{1}\otimes x_{r1}+\cdots+e_{n}\otimes x_{rn}$
the generic elements of $\widehat{\mathcal{A}}_{m}$, where $1\leqslant r\leqslant m$. In particular, we have
$g\bullet X_{r}=\left(\begin{array}[]{c}gx_{r1}\\ \vdots\\ gx_{rn}\\ \end{array}\right)\in\widehat{\mathcal{A}}_{m}.$
For $g\in G\leqslant{\rm GL}_{n}$ we write $[g]$ for the corresponding $n\times n$ matrix. By straightforward calculations we can see that
$g\bullet X_{r}=[g]^{-1}X_{r}.$ (2.1)
Denote by ${\rm alg}_{\mathbb{F}}\{X\}_{m}$ the subalgebra of $\widehat{\mathcal{A}}_{m}$ generated by the generic elements $X_{1},\ldots,X_{m}$ and the unity $1$. Any product of the generic elements is called a word of ${\rm alg}_{\mathbb{F}}\{X\}_{m}$. Since $G\leqslant\mathop{\rm Aut}(\mathcal{A})$, we have that
$g\bullet(FH)=(g\bullet F)(g\bullet H)$ (2.2)
for all $F,H$ from ${\rm alg}_{\mathbb{F}}\{X\}_{m}$.
2.2 Operator traces
For $a\in\mathcal{A}$ denote by $L_{a}:\mathcal{A}\to\mathcal{A}$ and $R_{a}:\mathcal{A}\to\mathcal{A}$ the operators of the left and right multiplication by $a$, respectively. Then define the left operator trace $\mathop{\rm tr}_{\rm L}:\mathcal{A}\to\mathbb{F}$ by $\mathop{\rm tr}_{\rm L}(a)=\mathop{\rm tr}(L_{a})$ and the right operator trace $\mathop{\rm tr}_{\rm R}:\mathcal{A}\to\mathbb{F}$ by $\mathop{\rm tr}_{\rm R}(a)=\mathop{\rm tr}(R_{a})$.
To expand these constructions, we denote by $\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ the absolutely free unital algebra in letters $\chi_{0},\chi_{1},\ldots,\chi_{m}$ and for $f\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$, ${\underline{a}}=(a_{0},\ldots,a_{m})\in\mathcal{A}^{m+1}$ define $f({\underline{a}})\in\mathcal{A}$ as the result of substitutions $\chi_{0}\to a_{0},\,\chi_{1}\to a_{1},\ldots,\chi_{m}\to a_{m}$ in $f$. For $h\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ denote by $L_{h},R_{h}:\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle\to\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ the operators of the left and right multiplication by $h$, respectively. As usually, the composition of maps $P_{1},P_{2}:\mathcal{A}\to\mathcal{A}$ is denoted by $P_{1}\circ P_{2}(a)=P_{1}(P_{2}(a))$, $a\in\mathcal{A}$. Similar notation we use for composition of maps $P_{1},P_{2}:\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle\to\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$. For a symbol $P\in\{L,R\}$, we write
  1. $\bullet$
    $P_{a}:\mathcal{A}\to\mathcal{A}$ for the operator of left or right multiplication by $a\in\mathcal{A}$;
  2. $\bullet$
    $P_{h}:\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle\to\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ for the operator of left or right multiplication by $h\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$.
The following definition of the operator trace generalizes definitions of the left and right operator traces.
Definition 1.
Assume that $m\geqslant 1$ and $h\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ is homogeneous of degree 1 in $\chi_{0}$, i.e., each monomial of $h$ contains $\chi_{0}$ exactly once. Then
  1. $\bullet$
    for every ${\underline{a}}\in\mathcal{A}^{m}$ define the linear operator $h(\,\cdot\,,{\underline{a}}):\mathcal{A}\to\mathcal{A}$ by the following equality: $h(\,\cdot\,,{\underline{a}})(b)=h(b,a_{1},\ldots,a_{m})$ for all $b\in\mathcal{A}$;
  2. $\bullet$
    define operator trace as follows: $\mathop{\rm tr}(h):\mathcal{A}^{m}\to\mathbb{F}$, where $\mathop{\rm tr}(h)({\underline{a}})=\mathop{\rm tr}(h(\,\cdot\,,{\underline{a}}))$ for all ${\underline{a}}\in\mathcal{A}^{m}$.
As an example, $\mathop{\rm tr}(\chi_{0})=n$, $\mathop{\rm tr}_{\rm L}=\mathop{\rm tr}(\chi_{1}\chi_{0})$, and $\mathop{\rm tr}_{\rm R}=\mathop{\rm tr}(\chi_{0}\chi_{1})$.
Remark 2.
Assume that $h\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ is homogeneous of degree 1 in $\chi_{0}$.
  1. (a)
    It is easy to see that $\mathop{\rm tr}(h)\in\mathbb{F}[\mathcal{A}^{m}]$.
  2. (b)
    It is well known that $\mathop{\rm tr}(h)$ is an invariant from $I_{m}(\mathcal{A})$ (for example, see Lemma 3.1 from [1]).
  3. (c)
    If $h$ has multidegree $(1,\delta_{1},\ldots,\delta_{m})\in\mathbb{N}^{m+1}$, then $\mathop{\rm tr}(h)$ is $\mathbb{N}^{m}$-homogeneous of multidegree $(\delta_{1},\ldots,\delta_{m})$.
Definition 3.
Denote by $\mathop{\rm tr}(\mathcal{A})_{m}$ the subalgebra of $I_{m}(\mathcal{A})$ generated by $1$ and operator traces $\mathop{\rm tr}(h)$ for all that $h\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ homogeneous of degree 1 in $\chi_{0}$.
3 Generators for algebras of invariants
3.1 Traces
Assume that the tableau of multiplication of the algebra $\mathcal{A}$ is $M=(M_{ij})_{1\leqslant i,j\leqslant n}$, i.e., $e_{i}e_{j}=M_{ij}$ in $\mathcal{A}$. Denote
$M_{ij}=\sum_{l=1}^{n}M_{ijl}e_{l}$
for some $M_{ijl}\in\mathbb{F}$. For every $1\leqslant i\leqslant n$ consider the following $n\times n$ matrices over $\mathbb{F}$:
$M_{\rm L}^{(i)}=(M_{ijl})_{1\leqslant l,j\leqslant n}\;\;\text{ and }\;\;M_{\rm R}^{(i)}=(M_{jil})_{1\leqslant l,j\leqslant n}.$
Proposition 4 (Proposition 3.2 of [1]).
Assume that $m\geqslant 1$ and a homogeneous monomial $h\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ of degree 1 in $\chi_{0}$ satisfies the equality $h=P^{1}_{\chi_{r_{1}}}\circ\cdots\circ P^{k}_{\chi_{r_{k}}}(\chi_{0})$ for some symbols $P^{1},\ldots,P^{k}\in\{L,R\}$, $1\leqslant r_{1},\ldots,r_{k}\leqslant m$ and $k>0$. Then
$\mathop{\rm tr}(h)=\sum_{1\leqslant i_{1},\ldots,i_{k}\leqslant n}\mathop{\rm tr}\!\Big{(}M_{P^{1}}^{(i_{1})}\cdots M_{P^{k}}^{(i_{k})}\Big{)}x_{r_{1},i_{1}}\cdots x_{r_{k},i_{k}}.$
Proposition 4 implies that for all $1\leqslant r\leqslant m$ we have
$\mathop{\rm tr}(\chi_{r}\chi_{0})=\sum_{i,j=1}^{n}x_{ri}M_{ijj}\;\;\text{ and }\;\;\mathop{\rm tr}(\chi_{0}\chi_{r})=\sum_{i,j=1}^{n}x_{ri}M_{jij}.$ (3.1)
For every $1\leqslant i,i^{\prime}\leqslant n$ consider the following $n\times n$ matrices over $\mathbb{F}$:
$M_{\rm L}^{(i,i^{\prime})}=\Big{(}\sum_{t=1}^{n}M_{ii^{\prime}t}M_{tjl}\Big{)}_{1\leqslant l,j\leqslant n}\;\;\text{ and }\;\;M_{\rm R}^{(i,i^{\prime})}=\Big{(}\sum_{t=1}^{n}M_{ii^{\prime}t}M_{jtl}\Big{)}_{1\leqslant l,j\leqslant n}.$
Proposition 5 (Proposition 3.3 of [1]).
Assume that $m\geqslant 1$ and a homogeneous monomial $h\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ of degree 1 in $\chi_{0}$ satisfies the equality $h=P^{1}_{h_{1}}\circ\cdots\circ P^{k}_{h_{k}}(\chi_{0})$ for some symbols $P^{1},\ldots,P^{k}\in\{L,R\}$ and monomials $h_{1},\ldots,h_{k}\in\mathbb{F}\langle\chi_{1},\ldots,\chi_{m}\rangle$ of degree 1 or 2, where $k>0$. For every $1\leqslant q\leqslant k$ denote
  1. $\bullet$
    ${\underline{i}}_{q}=i_{q}$ and $w(q,{\underline{i}}_{q})=x_{r_{q},i_{q}}$, in case $h_{q}=\chi_{r_{q}}$ for some $1\leqslant r_{q}\leqslant m$;
  2. $\bullet$
    ${\underline{i}}_{q}=\{i_{q},i^{\prime}_{q}\}$, $M_{P^{q}}^{({\underline{i}}_{q})}=M_{P^{q}}^{(i_{q},i^{\prime}_{q})}$ and $w(q,{\underline{i}}_{q})=x_{r_{q},i_{q}}x_{r^{\prime}_{q},i^{\prime}_{q}}$, in case $h_{q}=\chi_{r_{q}}\chi_{r^{\prime}_{q}}$ for some $1\leqslant r_{q},r^{\prime}_{q}\leqslant m$.
Then
$\mathop{\rm tr}(h)=\sum_{1\leqslant{\underline{i}}_{1},\ldots,{\underline{i}}_{k}\leqslant n}\mathop{\rm tr}\!\Big{(}M_{P^{1}}^{({\underline{i}}_{1})}\cdots M_{P^{k}}^{({\underline{i}}_{k})}\Big{)}w(1,{\underline{i}}_{1})\cdots w(k,{\underline{i}}_{k}),$
where the condition $1\leqslant\{i_{q},i^{\prime}_{q}\}\leqslant n$ stands for the condition $1\leqslant i_{q},i^{\prime}_{q}\leqslant n$.
3.2 Reduction to multilinear case
In this section we assume that $G\leqslant{\rm GL}_{n}$ and $\mathbb{F}=\mathbb{C}$ is the field of complex numbers.
A vector ${\underline{r}}=(r_{1},\ldots,r_{m})\in\mathbb{N}^{m}$ satisfying $r_{1}+\cdots+r_{m}=t$ is called a partition of $t\in\mathbb{N}$ in $m$ parts. Denote by ${\mathcal{P}}^{t}_{m}$ the set of all such partitions. Given a partition ${\underline{r}}\in{\mathcal{P}}^{t}_{m}$, we define a function ${\underline{r}}|\cdot|:\{1,\ldots,t\}\to\{1,\ldots,m\}$ by
${\underline{r}}|l|=j\text{ if and only if }r_{1}+\cdots+r_{j-1}+1\leqslant l\leqslant r_{1}+\cdots+r_{j},$
for all $1\leqslant l\leqslant t$. For a partition ${\underline{r}}\in{\mathcal{P}}^{t}_{m}$ we also define a homomorphism of $\mathbb{F}$-algebras as follows:
$\pi_{{\underline{r}}}:\mathbb{F}[\mathcal{A}^{t}]\to\mathbb{F}[\mathcal{A}^{m}]$ by $x_{l,i}\to x_{{\underline{r}}|l|,i}$ for all $1\leqslant l\leqslant t$ and $1\leqslant i\leqslant n$.
The following proposition and lemma are well known and can easily be proven.
Proposition 6.
The algebra of invariants $\mathbb{F}[\mathcal{A}^{m}]^{G}$ is generated by
$\{\pi_{{\underline{r}}}(f)\,|\,f\in\mathbb{F}[\mathcal{A}^{t}]^{G}\text{ is multilinear},\;{\underline{r}}\in{\mathcal{P}}^{t}_{m},\;t\geqslant m\}.$
Lemma 7.
If the group $G$ is diagonal, i.e., all elements of $G\leqslant{\rm GL}_{n}$ are diagonal matrices, then the algebra of invariants $\mathbb{F}[\mathcal{A}^{m}]^{G}$ is generated by some monomials from $\mathbb{F}[\mathcal{A}^{m}]$.
The following lemma follows immediately from a partial case of Theorem 7.6 from [1] for the algebra ${\mathbf{A}}_{2}$.
Lemma 8.
Assume that $\mathcal{V}$ is a two-dimensional vector space with a basis $\{e_{1},e_{2}\}$ and let $G\leqslant{\rm GL}_{2}$ be the group of all matrices
$\left(\begin{array}[]{cc}1&\alpha\\ 0&1\\ \end{array}\right),$
[…]
$1,x_{12},\ldots,x_{m2},\;x_{r1}x_{s2}-x_{r2}x_{s1}\;\;(1\leqslant r<s\leqslant m).$
4 Automorphisms of $3$-dimensional Leibniz algebras
We use notations for the tableau of multiplication of an algebra from Section 3.1.
Theorem 9 (Rikhsiboev, Rakhimov [34], Theorem 2).
Each algebra $\mathcal{L}\in{\bf L}_{3}$ is isomorphic to one and only one algebra from the following list, where algebras are given by their tableaux of multiplication with respect to some basis $\{e_{1},e_{2},e_{3}\}$:
$\mathcal{L}_{1}$ :::: $\left(\begin{array}[]{ccc}0&0&-2e_{1}\\ 0&e_{1}&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{2}^{\lambda}$ :::: $\left(\begin{array}[]{ccc}0&0&\lambda e_{1}\\ 0&0&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{3}$ :::: $\left(\begin{array}[]{ccc}0&0&0\\ 0&0&-e_{2}\\ 0&e_{2}&e_{1}\\ \end{array}\right),$
$\mathcal{L}_{4}^{\lambda}$ :::: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&e_{1}\\ 0&0&\lambda e_{1}\\ \end{array}\right),$ $\mathcal{L}_{5}$ :::: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{6}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&e_{1}\\ 0&0&0\\ \end{array}\right),$
$\mathcal{L}_{7}^{\lambda}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&\lambda e_{1}+e_{2}\\ 0&0&0\\ \end{array}\right),$ $\mathcal{L}_{8}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{9}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{1}+e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$
where $\lambda\in\mathbb{F}$.
Theorem 10.
The groups of automorphisms of algebras from ${\bf L}_{3}$ are the following set of matrices:
$\mathcal{L}_{1}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{5}^{2}&\alpha_{5}\alpha_{6}&\alpha_{6}^{2}/2\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
$\mathcal{L}_{2}^{\lambda\neq 0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&0\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1},\alpha_{5}\neq 0;$
$\mathcal{L}_{2}^{0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1},\alpha_{5}\neq 0;$
$\mathcal{L}_{3}$ :::: $g=\left(\begin{array}[]{ccc}1&0&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
$\mathcal{L}_{4}^{\lambda\neq 0}$ :::: $g_{1}=\left(\begin{array}[]{ccc}\frac{\alpha_{6}^{2}+\alpha_{6}\alpha_{9}+\lambda\alpha_{9}^{2}}{\lambda}&\alpha_{2}&\alpha_{3}\\ 0&\frac{\alpha_{6}}{\lambda}+\alpha_{9}&\alpha_{6}\\ 0&-\frac{\alpha_{6}}{\lambda}&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{6}\neq 0,\;\det(g_{1})\neq 0,$
$g_{2}=\left(\begin{array}[]{ccc}\alpha_{5}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&0\\ 0&0&\alpha_{5}\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
$\mathcal{L}_{4}^{0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{5}(\alpha_{5}+\alpha_{8})&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&0\\ 0&\alpha_{8}&\alpha_{5}+\alpha_{8}\\ \end{array}\right)\text{ for }\alpha_{5}+\alpha_{8}\neq 0;$
$\mathcal{L}_{5}$ :::: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{5}^{2}+\alpha_{6}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&-\alpha_{6}&\alpha_{5}\\ \end{array}\right)\text{ for }a_{5}\neq 0,\;\alpha_{5}^{2}+\alpha_{6}^{2}\neq 0,$
$g_{2}=\left(\begin{array}[]{ccc}\alpha_{5}^{2}+\alpha_{6}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&\alpha_{6}&-\alpha_{5}\\ \end{array}\right)\text{ for }a_{5}\neq 0,\;\alpha_{5}^{2}+\alpha_{6}^{2}\neq 0,$
$g_{3}=\left(\begin{array}[]{ccc}\alpha_{6}^{2}&\alpha_{2}&\alpha_{3}\\ 0&0&\alpha_{6}\\ 0&\pm\alpha_{6}&0\\ \end{array}\right)\text{ for }\alpha_{6}\neq 0;$
$\mathcal{L}_{6}$ :::: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{1}&\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}\neq\alpha_{4}^{2};$
$g_{2}=\left(\begin{array}[]{ccc}\alpha_{1}&-\alpha_{4}&0\\ \alpha_{4}&-\alpha_{1}&0\\ 0&0&-1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}\neq\alpha_{4}^{2};$
$\mathcal{L}_{7}^{\lambda\neq 0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&\lambda\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}+\alpha_{1}\alpha_{4}-\lambda\alpha_{4}^{2}\neq 0;$
$\mathcal{L}_{7}^{0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&\alpha_{3}\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&-\alpha_{3}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}(\alpha_{1}+\alpha_{4})\neq 0;$
$\mathcal{L}_{8}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{9}^{2}&0&\alpha_{3}\\ \alpha_{3}\alpha_{9}&\alpha_{9}^{3}&\alpha_{6}\\ 0&0&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{9}\neq 0;$
$\mathcal{L}_{9}$ :::: $g=\left(\begin{array}[]{ccc}1+\alpha_{3}&0&\alpha_{3}\\ \alpha_{3}&1&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{3}\neq-1.$
Here $\alpha_{1},\ldots,\alpha_{9}\in\mathbb{F}$.
Proof.
Given $\mathcal{L}\in{\bf L}_{3}$, consider an invertible linear map $g:\mathcal{L}\to\mathcal{L}$, given by a matrix
$g=\left(\begin{array}[]{ccc}\alpha_{1}&\alpha_{2}&\alpha_{3}\\ \alpha_{4}&\alpha_{5}&\alpha_{6}\\ \alpha_{7}&\alpha_{8}&\alpha_{9}\\ \end{array}\right)$
with respect to the basis $\{e_{1},e_{2},e_{3}\}$ of $\mathcal{L}$ from Theorem 9, where $\alpha_{1},\ldots,\alpha_{9}\in\mathbb{F}$. Then $g$ is an automorphism of $\mathcal{L}$ if and only if for each $1\leqslant i,j\leqslant 3$ we have $g_{ij}:=g(e_{i}e_{j})-g(e_{i})g(e_{j})=0$. It is straightforward to verify that matrices $g,g_{i}$ from the formulation of this theorem are automorphisms. So, we assume that $g$ is an automorphism.
  1. 1.
    Assume $\mathcal{L}=\mathcal{L}_{1}$. Consequently considering $g_{13}=0$, $g_{11}=0$, $g_{23}=0$ we obtain $\alpha_{7}=0$, $\alpha_{4}=0$, $\alpha_{8}=0$, respectively. Since $\det(g)=\alpha_{1}\alpha_{5}\alpha_{9}$, we have $\alpha_{1}\neq 0$ and consequently considering $g_{13}=0$, $g_{22}=0$, $g_{23}=0$, $g_{33}=0$ we obtain $\alpha_{9}=1$, $\alpha_{1}=\alpha_{5}^{2}$, $\alpha_{2}=\alpha_{5}\alpha_{6}$, $\alpha_{3}=\alpha_{6}^{2}/2$, respectively. The required statement is proven.
  2. 2.
    Assume $\mathcal{L}=\mathcal{L}_{2}^{\lambda}$ for some $\lambda\in\mathbb{F}$.
    1. (a)
      Let $\lambda\neq 0$. Consequently considering $g_{13}=0$, $g_{23}=0$, $g_{32}=0$ we obtain $\alpha_{7}=0$, $\alpha_{8}=0$, $\alpha_{2}=0$, respectively. Since $\det(g)=\alpha_{1}\alpha_{5}\alpha_{9}\neq 0$, we consequently consider $g_{31}=0$, $g_{33}=0$, $g_{13}=0$ to obtain $\alpha_{4}=0$, $\alpha_{3}=0$, $\alpha_{9}=1$, respectively.
    2. (b)
      Let $\lambda=0$. Since $g_{32}=0$, we have $\alpha_{2}=0$ and $\alpha_{8}=0$. Since $\det(g)=-\alpha_{3}\alpha_{5}\alpha_{7}+\alpha_{1}\alpha_{5}\alpha_{9}$, we have $\alpha_{5}\neq 0$. Since $g_{12}=0$, we have $\alpha_{7}=0$. Thus $\det(g)=\alpha_{1}\alpha_{5}\alpha_{9}\neq 0$ and we consequently consider $g_{13}=0$, $g_{23}=0$ to obtain $\alpha_{4}=0$, $\alpha_{9}=1$, respectively. The required statement is proven.
  3. 3.
    Assume $\mathcal{L}=\mathcal{L}_{3}$. Consequently considering $g_{11}=0$, $g_{22}=0$, $g_{23}=0$, we obtain $\alpha_{7}=0$, $\alpha_{8}=0$, $\alpha_{2}=0$, respectively. Since $\det(g)=\alpha_{1}\alpha_{5}\alpha_{9}\neq 0$, we consequently consider $g_{13}=0$, $g_{23}=0$, $g_{33}=0$ to obtain $\alpha_{4}=0$, $\alpha_{9}=1$, $\alpha_{1}=1$, respectively. The required statement is proven.
  4. 4.
    Assume $\mathcal{L}=\mathcal{L}_{4}^{\lambda}$ for some $\lambda\in\mathbb{F}$. Since $g_{22}=0$, we have $\alpha_{4}=0$ and $\alpha_{7}=0$. Considering equality $g_{22}=0$ once again, we can see that $\alpha_{1}=\alpha_{5}^{2}+\alpha_{5}\alpha_{8}+\lambda\alpha_{8}^{2}$.
    1. (a)
      Let $\lambda\neq 0$.
      1. i.
        Assume $\alpha_{6}\neq 0$. Then equality $g_{32}=0$ implies that $\alpha_{5}=-\alpha_{8}(\alpha_{6}+\lambda\alpha_{9})/\alpha_{6}$. Since $\det(g)\neq 0$, we have $\alpha_{6}^{2}+\alpha_{6}\alpha_{9}+\lambda\alpha_{9}^{2}\neq 0$ and $\alpha_{8}\neq 0$. Thus $g_{23}=0$ implies $\alpha_{8}=-\alpha_{6}/\lambda$.
      2. ii.
        Assume $\alpha_{6}=0$. Since $\det(g)\neq 0$, we have $\alpha_{5}\neq 0$ and $\alpha_{9}\neq 0$. Consequently considering $g_{32}=0$ and $g_{23}=0$ we obtain that $\alpha_{8}=0$ and $\alpha_{9}=\alpha_{5}$, respectively.
    2. (b)
      Let $\lambda=0$. Since $\det(g)\neq 0$, we have $\alpha_{5}+\alpha_{8}\neq 0$ and $\alpha_{5}\neq 0$. Consequently considering $g_{32}=0$ and $g_{23}=0$ we obtain $\alpha_{6}=0$ and $\alpha_{9}=\alpha_{5}+\alpha_{8}$, respectively. The required statement is proven.
  5. 5.
    Assume $\mathcal{L}=\mathcal{L}_{5}$. Equality $g_{22}=0$ implies $\alpha_{4}=0$, $\alpha_{7}=0$, and then equality $g_{33}=0$ implies $\alpha_{1}=\alpha_{6}^{2}+\alpha_{9}^{2}$.
    1. (a)
      Let $\alpha_{9}\neq 0$. Equality $g_{23}=0$ implies that $\alpha_{8}=-\alpha_{5}\alpha_{6}/\alpha_{9}$. Since $\det(g)\neq 0$, we have $\alpha_{6}^{2}+\alpha_{9}^{2}\neq 0$ and $g_{22}=0$ implies that $\alpha_{9}=\pm\alpha_{5}$.
    2. (b)
      Let $\alpha_{9}=0$. Since $\det(g)=-\alpha_{6}^{3}\alpha_{8}\neq 0$, consequently considering $g_{23}=0$, $g_{22}=0$ we obtain $\alpha_{5}=0$ and $\alpha_{8}=\pm\alpha_{6}$, respectively. The required statement is proven.
  6. 6.
    Assume $\mathcal{L}=\mathcal{L}_{6}$. Consequently considering $g_{13}=0$, $g_{23}=0$ we obtain that $\alpha_{8}=0$ and $\alpha_{7}=0$, respectively. Considering equality $g_{13}=0$ once again, we can see that $\alpha_{5}=\alpha_{1}\alpha_{9}$ and $\alpha_{2}=\alpha_{4}\alpha_{9}$. Since $\det(g)=(\alpha_{1}^{2}-\alpha_{4}^{2})\alpha_{9}^{2}\neq 0$, equality $g_{33}=0$ implies $\alpha_{3}=\alpha_{6}=0$.
    If $\alpha_{9}\neq\pm 1$, then $g_{23}=0$ implies $\alpha_{1}=\alpha_{4}=0$; hence, $\det(g)=0$; a contradiction. Therefore, $\alpha_{9}=\pm 1$. The required statement is proven.
  7. 7.
    Assume $\mathcal{L}=\mathcal{L}_{7}^{\lambda}$ for some $\lambda\in\mathbb{F}$. Considering $g_{13}=0$ we obtain $\alpha_{8}=0$, $\alpha_{2}=\lambda\alpha_{4}\alpha_{9}$ and $\alpha_{5}=(\alpha_{1}+\alpha_{4})\alpha_{9}$. Note that $\alpha_{9}\neq 0$, since $\det(g)\neq 0$. Hence, it follows from $g_{33}=0$ that $\alpha_{6}=-\alpha_{3}$.
    1. (a)
      Let $\lambda\neq 0$. Consequently considering $g_{23}=0$ and $g_{33}=0$, we obtain $\alpha_{7}=0$ and $\alpha_{3}=0$, respectively.
      1. i.
        Assume $\alpha_{9}=-1$. Since the characteristic of the field is zero, $g_{23}=0$ implies $\alpha_{4}=0$ and $\alpha_{1}=0$. Thus, $\det(g)=0$; a contradiction.
      2. ii.
        Assume $\alpha_{9}\neq\pm 1$. Then $g_{23}=0$ implies $\alpha_{1}=-\alpha_{4}\alpha_{9}/(1+\alpha_{9})$. Considering $g_{23}=0$ once again we obtain $\alpha_{4}(\alpha_{9}+\lambda(1+\alpha_{9})^{2})=0$. Thus $\det(g)=0$; a contradiction.
      Therefore, $\alpha_{9}=1$.
    2. (b)
      Let $\lambda=0$. Note that $\det(g)=(\alpha_{1}+\alpha_{4})\alpha_{9}(-\alpha_{3}\alpha_{7}+\alpha_{1}\alpha_{9})$. Since $(\alpha_{1}+\alpha_{4})\alpha_{9}\neq 0$, consequently considering $g_{11}=0$ and $g_{23}=0$ we can see that $\alpha_{7}=0$ and $\alpha_{9}=1$, respectively. The required statement is proven.
  8. 8.
    Assume $\mathcal{L}=\mathcal{L}_{8}$ for some $\lambda\in\mathbb{F}$. Equality $g_{13}=0$ implies that $\alpha_{8}=0$, $\alpha_{5}=\alpha_{1}\alpha_{9}$, $\alpha_{2}=\alpha_{7}\alpha_{9}$. Hence, it follows from equality $g_{33}=0$ that $\alpha_{7}=0$, $\alpha_{4}=\alpha_{3}\alpha_{9}$, $\alpha_{1}=\alpha_{9}^{2}$. The required statement is proven.
  9. 9.
    Assume $\mathcal{L}=\mathcal{L}_{9}$. Equality $g_{33}=0$ implies that $\alpha_{4}=\alpha_{3}\alpha_{9}$. Applying equality $g_{33}=0$ once again, we can see that $\alpha_{7}=0$ and $\alpha_{1}=\alpha_{9}(\alpha_{3}+\alpha_{9})$. Then equality $g_{13}=0$ implies $\alpha_{8}=0$. Thus $\alpha_{9}\neq 0$, since $\det(g)=\alpha_{9}^{2}(-\alpha_{2}\alpha_{3}+\alpha_{5}(\alpha_{3}+\alpha_{9}))$. Consequently considering $g_{23}=0$ and $g_{13}=0$ we obtain that $\alpha_{2}=0$ and $\alpha_{5}=-\alpha_{3}\alpha_{9}+\alpha_{3}\alpha_{9}^{2}+\alpha_{9}^{3}$, respectively.
    In case $\alpha_{9}=-\alpha_{3}$ we have $\det(g)=0$; a contradiction. Therefore, $\alpha_{9}\neq-\alpha_{3}$ and $g_{13}=0$ implies that $\alpha_{9}=1$. The required statement is proven.
Corollary 11.
Given $\mathcal{L}\in{\bf L}_{3}$, the group $\mathop{\rm Aut}(\mathcal{L})$ has dimension $2$, $3$ or $4$ as an affine variety, where each of the possibilities holds for some algebra $\mathcal{L}\in{\bf L}_{3}$.
An automorphism $\delta$ of an algebra $\mathcal{A}$ is called diagonal if there exists some basis $\{e_{1},\ldots,e_{n}\}$ of $\mathcal{A}$ such that $\varphi(e_{i})=\alpha_{i}e_{i}$ for some $\alpha_{i}\in\mathbb{F}$. For short, we write a diagonal $n\times n$ matrix with diagonal elements $\alpha_{1},\ldots,\alpha_{n}\in\mathbb{F}$ as ${\rm diag}\{\alpha_{1},\ldots,\alpha_{n}\}$. The importance of the diagonal automorphisms for invariant theory follows from the next remark.
Remark 12.
[…]
Theorem 10 implies the following lemma.
Lemma 13.
Here is a list of some diagonal automorphisms $\delta$:
$\mathcal{L}_{1}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha,1\},$ $\mathcal{L}_{2}^{\lambda}$ :::: $\delta={\rm diag}\{\alpha,\beta,1\},$
$\mathcal{L}_{3}$ :::: $\delta={\rm diag}\{1,\alpha,1\},$ $\mathcal{L}_{4}^{\lambda}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha,\alpha\},$
$\mathcal{L}_{5}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha,\pm\alpha\},$ $\mathcal{L}_{6}$ :::: $\delta={\rm diag}\{\alpha,\alpha,1\},$
$\mathcal{L}_{7}^{\lambda}$ :::: $\delta={\rm diag}\{\alpha,\alpha,1\},$ $\mathcal{L}_{8}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha^{3},\alpha\},$
for all $\lambda\in\mathbb{F}$ and all non-zero $\alpha,\beta\in\mathbb{F}$.
5 Generators for algebras of invariants
In this section we describe the algebra of polynomial invariants $I_{m}(\mathcal{L})$ for any $\mathcal{L}\in{\bf L}_{3}$. For short, we denote $x_{r}:=x_{r1}$, $y_{r}:=x_{r2}$, $z_{r}:=x_{r3}$ for all $1\leqslant r\leqslant m$. Given ${\underline{r}}=(r_{1},\ldots,r_{k})\in\{1,\ldots,m\}^{k}$, we write $x_{{\underline{r}}}=x_{r_{1}}\cdots x_{r_{k}}$, $y_{{\underline{r}}}=y_{r_{1}}\cdots y_{r_{k}}$, $z_{{\underline{r}}}=z_{r_{1}}\cdots z_{r_{k}}$.
Note that by formula (2.1) an automorphism $g=(g_{ij})_{1\leqslant i,j\leqslant 3}$ of $\mathop{\rm Aut}(\mathcal{L})\leqslant{\rm GL}_{3}$ acts on $\mathbb{F}[\mathcal{L}^{m}]$ as follows:
$\begin{array}[]{clc}g^{-1}x_{r}&=&g_{11}x_{r}+g_{12}y_{r}+g_{13}z_{r},\\ g^{-1}y_{r}&=&g_{21}x_{r}+g_{22}y_{r}+g_{23}z_{r},\\ g^{-1}z_{r}&=&g_{31}x_{r}+g_{32}y_{r}+g_{33}z_{r}\end{array}$ (5.1)
for all $1\leqslant r\leqslant m$.
Theorem 14.
For every $\lambda\in\mathbb{F}$ and $m>0$ we have the following equalities for algebras of invariants:
  1. $\bullet$
    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{\lambda})$, $I_{m}(\mathcal{L}_{7})$ are all equal to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
  2. $\bullet$
    $I_{m}(\mathcal{L}_{3})$ is generated by $1,z_{1},\ldots,z_{m}$ and $x_{r}z_{s}-z_{r}x_{s}$ $(1\leqslant r<s\leqslant m$),
  3. $\bullet$
    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})=\mathbb{F}$,
  4. $\bullet$
    $I_{m}(\mathcal{L}_{6})$ is generated by $1,z_{r}z_{s}$ for all $1\leqslant i\leqslant j\leqslant m$,
  5. $\bullet$
    $I_{m}(\mathcal{L}_{9})$ is generated by $1,z_{1},\ldots,z_{m}$ and $(x_{r}-y_{r})z_{s}-z_{r}(x_{s}-y_{s})$ for all $1\leqslant r<s\leqslant m$.
Proof.
Lemma 13 together with Remark 12 implies that for every $\lambda\in\mathbb{F}$ and $m>0$ we have the following inclusions and equalities for algebras of invariants:
  1. $\bullet$
    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{\lambda})$, $I_{m}(\mathcal{L}_{6})$, $I_{m}(\mathcal{L}_{7})$ belong to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
  2. $\bullet$
    $I_{m}(\mathcal{L}_{3})\subset\mathbb{F}[x_{1},\ldots,x_{m},z_{1},\ldots,z_{m}]$,
  3. $\bullet$
    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})=\mathbb{F}$.
Theorem 10 implies that $z_{1},\ldots,z_{m}\in I_{m}(\mathcal{L})$ for every $\mathcal{L}$ from the following list: $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{7}$, $\mathcal{L}_{9}$. Thus, the statement of the theorem holds for algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{4}^{\lambda}$, $\mathcal{L}_{5}$, $\mathcal{L}_{7}$, $\mathcal{L}_{8}$.
Consider the algebra $\mathcal{L}_{3}$. The subspace $V=\mathbb{F}\text{-span}\{e_{1},e_{3}\}$ is invariant with respect to $\mathop{\rm Aut}(\mathcal{L}_{3})$ and the restriction of $\mathop{\rm Aut}(\mathcal{L}_{3})$ to $V$ is the set of all matrices
$\left(\begin{array}[]{cc}1&\alpha\\ 0&1\\ \end{array}\right)$
for all $\alpha\in\mathbb{F}$. Then applying Lemma 8 we prove the statement about $\mathcal{L}_{3}$.
Consider the algebra $\mathcal{L}_{6}$. Note that $z_{r}z_{s}\in I_{m}(\mathcal{L}_{6})$ for all $1\leqslant r\leqslant s\leqslant m$ by Theorem 10. Then a monomial in $z_{1},\ldots,z_{m}$ lies in $I_{m}(\mathcal{L}_{6})$ if and only if it has even degree. Restricting the group $\mathop{\rm Aut}(\mathcal{L}_{6})$ to $\mathbb{F}\text{-span}\{e_{3}\}$ and applying Lemma 7 we conclude the proof about $\mathcal{L}_{6}$.
Consider the algebra $\mathcal{L}_{9}$. Using formlas (5.1) we can see that $(x_{r}-y_{r})z_{s}-z_{r}(x_{s}-y_{s})$ lies in $I_{m}(\mathcal{L}_{9})$. For every $\alpha\in\mathbb{F}\backslash\{-1\}$ consider an automorphism
$g=\left(\begin{array}[]{ccc}1+\alpha&0&\alpha\\ \alpha&1&0\\ 0&0&1\\ \end{array}\right)$
of $\mathcal{L}_{9}$ (see Theorem 10). For $e_{1}^{\prime}=e_{1}-e_{2}$ we have that
$ge_{1}^{\prime}=e_{1}^{\prime}+\alpha e_{3},\;\;ge_{2}=\alpha e^{\prime}_{1}+(\alpha+1)e_{2},\;\;ge_{3}=e_{3}.$ (5.2)
[…]
For each $m\geqslant 0$, denote by $\Omega_{m}$ the set of all monomials $w=w_{1}\cdots w_{m}\in\mathbb{F}[\mathcal{L}^{m}]$ of multidegree $1^{m}$ for some $w_{r}\in\{x^{\prime}_{r},y_{r},z_{r}\}$, where $1\leqslant r\leqslant m$. Given $w\in\Omega_{m}$, denote
$\begin{array}[]{rcl}\deg_{x^{\prime}}(w)&=&\deg_{x^{\prime}_{1}}(w)+\cdots+\deg_{x_{m}}(w),\\ \deg_{y}(w)&=&\deg_{y_{1}}(w)+\cdots+\deg_{y_{m}}(w),\\ \deg_{z}(w)&=&\deg_{z_{1}}(w)+\cdots+\deg_{z_{m}}(w).\\ \end{array}$
As an example, for $m=3$ we have $\deg_{y}(z_{1}x^{\prime}_{2}y_{3})=1$ and $\deg_{y}(x^{\prime}_{1}y_{2}y_{3})=2$. Introduce a complete order on $\Omega_{m}$ as follows: for $v=v_{1}\cdots v_{m}$ and $w=w_{1}\cdots w_{m}$ from $\Omega_{m}$ we have $v<w$ if and only if
  1. (a)
    $\deg_{y}(v)<\deg_{y}(w)$, or
  2. (b)
    $\deg_{y}(v)=\deg_{y}(w)$ and $\deg_{x}(v)<\deg_{x}(w)$, or
  3. (c)
    $\deg_{y}(v)=\deg_{y}(w)$, $\deg_{x}(v)=\deg_{x}(w)$, and there exists $1\leqslant r\leqslant m$ with
    1. $\bullet$
      $v_{1}=w_{1},\ldots,v_{r-1}=w_{r-1}$,
    2. $\bullet$
      $(v_{r},w_{r})$ is one of the following pairs: $(x^{\prime}_{r},y_{r})$, $(z_{r},y_{r})$, $(z_{r},x^{\prime}_{r})$.
As an example, for $m=3$ we have
$x_{1}^{\prime}x_{2}^{\prime}y_{3}<z_{1}y_{2}y_{3},\;\;\;\;x_{1}^{\prime}x_{2}^{\prime}y_{3}<z_{1}x^{\prime}_{2}y_{3},\;\;\text{ and }\;\;y_{1}z_{2}x^{\prime}_{3}<y_{1}x^{\prime}_{2}y_{3}.$
For a multilinear $f=\sum_{j}\beta_{j}w_{j}\in\mathbb{F}[\mathcal{L}^{m}]$ with non-zero $\beta_{j}\in\mathbb{F}$ and $w_{j}\in\Omega_{m}$ we say that $w_{j_{0}}$ is the highest term for $f$, if there is no $j\neq j_{0}$ with $w_{j}>w_{j_{0}}$. Note that the highest term for $f$ always exists and is unique.
Consider a multilinear $f=\beta w+\sum_{j}\beta_{j}w_{j}\in I_{m}(\mathcal{L})$ with non-zero $\beta,\beta_{j}\in\mathbb{F}$ and $w,w_{j}\in\Omega_{m}$, where $w$ is the highest term for $f$. Without loss of generality we can assume that
$w=x^{\prime}_{1}\cdots x^{\prime}_{r}y_{r+1}\cdots y_{s}z_{s+1}\cdots z_{m},$
where $0\leqslant r\leqslant s\leqslant m$. Formulas (5.1) and (5.2) and imply that
$g^{-1}f=\beta(x_{1}^{\prime}+\alpha z_{1})\cdots(x_{r}^{\prime}+\alpha z_{r})\big{(}\alpha x^{\prime}_{r+1}+(\alpha+1)y_{r+1}\big{)}\cdots\big{(}\alpha x^{\prime}_{s}+(\alpha+1)y_{s}\big{)}z_{s+1}\cdots z_{m}+\sum_{j}\beta_{j}g^{-1}w_{j}.$
[…]
Note that the subspace $V=\mathbb{F}\text{-span}\{e^{\prime}_{1},e_{3}\}$ is invariant with respect to $g$ and the restriction of $g$ to $V$ is
$\left(\begin{array}[]{cc}1&\alpha\\ 0&1\\ \end{array}\right)$
[…]
6 Trace invariants
Assume than $n=3$. Writing down $M_{ij}$ as $M_{ij}=(M_{ij1},M_{ij2},M_{ij3})$, we can see that formula (3.1) implies that
$\mathop{\rm tr}(\chi_{r}\chi_{0})=\left(\sum_{i=1}^{3}\alpha_{i}\right)x_{r}+\left(\sum_{i=1}^{3}\beta_{i}\right)y_{r}+\left(\sum_{i=1}^{3}\gamma_{i}\right)z_{r}\;\text{ for }\;M=\left(\begin{array}[]{ccc}(\alpha_{1},\ast,\ast)&(\ast,\alpha_{2},\ast)&(\ast,\ast,\alpha_{3})\\ (\beta_{1},\ast,\ast)&(\ast,\beta_{2},\ast)&(\ast,\ast,\beta_{3})\\ (\gamma_{1},\ast,\ast)&(\ast,\gamma_{2},\ast)&(\ast,\ast,\gamma_{3})\\ \end{array}\right),$ (6.1)
$\mathop{\rm tr}(\chi_{0}\chi_{r})=\left(\sum_{i=1}^{3}\alpha_{i}\right)x_{r}+\left(\sum_{i=1}^{3}\beta_{i}\right)y_{r}+\left(\sum_{i=1}^{3}\gamma_{i}\right)z_{r}\;\text{ for }\;M=\left(\begin{array}[]{ccc}(\alpha_{1},\ast,\ast)&(\beta_{1},\ast,\ast)&(\gamma_{1},\ast,\ast)\\ (\ast,\alpha_{2},\ast)&(\ast,\beta_{2},\ast)&(\ast,\gamma_{2},\ast)\\ (\ast,\ast,\alpha_{3})&(\ast,\ast,\beta_{3})&(\ast,\ast,\gamma_{3})\\ \end{array}\right).$ (6.2)
Proposition 15.
For every $\lambda\in\mathbb{F}$, $m>0$ and $1\leqslant r,s\leqslant m$ we have the following trace formulas:
$\mathcal{L}_{1}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=z_{r},$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=-3z_{r},$
$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=z_{r}z_{s},$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=-z_{r}z_{s},$
$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=-z_{r}z_{s},$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=5z_{r}z_{s}.$
$\mathcal{L}_{2}^{\lambda}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=z_{r},$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=(\lambda-1)z_{r},$
$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=z_{r}z_{s},$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=-z_{r}z_{s},$
$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=-z_{r}z_{s},$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=(1+\lambda^{2})z_{r}z_{s}.$
$\mathcal{L}_{3}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=z_{r},$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=-z_{r},$
$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=z_{r}z_{s},$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=-z_{r}z_{s},$
$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=-z_{r}z_{s},$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=z_{r}z_{s},$
$\mathop{\rm tr}((\chi_{r}\chi_{s})\chi_{0}))=0,$ $\mathop{\rm tr}(\chi_{0}(\chi_{r}\chi_{s}))=0.$
$\mathcal{L}_{4}^{\lambda}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=0,$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=0,$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=0.$
$\mathcal{L}_{5}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=0,$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=0,$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=0.$
$\mathcal{L}_{6}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=0,$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=0,$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=2z_{r}z_{s}.$
$\mathcal{L}_{7}^{\lambda}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=z_{r},$
$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=0,$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=0,$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=(1+2\lambda)z_{r}z_{s}.$
$\mathcal{L}_{8}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=0,$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=0,$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=0.$
$\mathcal{L}_{9}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=z_{r},$
$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=0,$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=0,$
$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=0,$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=z_{r}z_{s},$
$\mathop{\rm tr}((\chi_{r}\chi_{s})\chi_{0}))=0,$ $\mathop{\rm tr}(\chi_{0}(\chi_{r}\chi_{s}))=0.$
Proof.
We take tableaux of multiplication for Leibniz algebras from Theorem 9. To calculate traces $\mathop{\rm tr}(\chi_{r}\chi_{0})$ and $\mathop{\rm tr}(\chi_{0}\chi_{r})$ we apply equalities (6.1) and (6.2), respectively. To calculate the rest of the traces we apply Proposition 4. ∎
Corollary 16.
Artin–Procesi–Iltyakov Equality does not hold for $\mathcal{L}\in{\bf L}_{3}$ if and only if $\mathcal{L}$ is isomorphic to $\mathcal{L}_{3}$ or $\mathcal{L}_{9}$.
Proof.
We apply Proposition 4, Theorem 14, part (c) of Remark 2 and the fact that the algebra $I_{m}(\mathcal{L})$ has $\mathbb{N}^{m}$-grading by multidegrees. ∎
Proposition 17.
We have
  1. $\bullet$
    ${\rm ncl}(\mathcal{L}_{4}^{\lambda})={\rm ncl}(\mathcal{L}_{5})=3$ for all $\lambda\in\mathbb{F}$;
  2. $\bullet$
    ${\rm ncl}(\mathcal{L}_{8})=4$;
  3. $\bullet$
    the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$, $\mathcal{L}_{9}$ are not nilpotent for all $\lambda\in\mathbb{F}$.
Proof.
For the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$ the elements $(\cdots((e_{2}e_{3})e_{3})\cdots)e_{3}$ are non-zero. Similarly, $(\cdots((e_{1}e_{3})e_{3})\cdots)e_{3}\neq 0$ in $\mathcal{L}_{9}$. Since $(e_{3}e_{3})e_{3}=e_{2}\neq 0$ in $\mathcal{L}_{8}$, we can see that ${\rm ncl}(\mathcal{L}_{8})=4$. The cases of $\mathcal{L}_{4}^{\lambda}$ and $\mathcal{L}_{5}$ are trivial. ∎
Corollary 18.
For an algebra $\mathcal{L}\in{\bf L}_{3}$, the following four conditions are equivalent:
  1. (a)
    the algebra of invariants $I_{m}(\mathcal{L})$ is $\mathbb{F}$ for all $m\geqslant 1$;
  2. (b)
    the algebra of invariants $I_{1}(\mathcal{L})$ is $\mathbb{F}$;
  3. (c)
    the algebra $\mathcal{L}$ is nilpotent;
  4. (d)
    $\mathop{\rm tr}(\chi_{0}\chi_{1})=\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=0$ for $\mathcal{L}$.
Proof.
The claim of the corollary follows from Proposition 17, Theorem 14 and Proposition 15. ∎
Corollary 19.
Two non-nilpotent algebras from ${\bf L}_{3}$ can be distinguished by means of the traces of degrees $\leqslant 2$ and the dimensions of its groups of automorphisms. That is, given $\mathcal{L}\in{\bf L}_{3}$, denote $d=\dim\mathop{\rm Aut}(\mathcal{L})$. Then
$\mathcal{L}\simeq\mathcal{L}_{1}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=2.$
$\mathcal{L}\simeq\mathcal{L}_{2}^{\lambda\neq 0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=(\lambda-1)\mathop{\rm tr}(\chi_{1}\chi_{0}).$
$\mathcal{L}\simeq\mathcal{L}_{2}^{0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=4.$
$\mathcal{L}\simeq\mathcal{L}_{3}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=-\mathop{\rm tr}(\chi_{1}\chi_{0}).$
$\mathcal{L}\simeq\mathcal{L}_{6}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=0.$
$\mathcal{L}\simeq\mathcal{L}_{7}^{\lambda\neq 0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=(1+2\lambda)\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$
$\mathcal{L}\simeq\mathcal{L}_{7}^{0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=3.$
$\mathcal{L}\simeq\mathcal{L}_{9}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$

### v2

@@
-Polynomial invariants for $3$-dimensional Leibniz algebras 111 The work is supported by the FCT 2023.08031.CEECIND, UIDB/00212/2020, UIDP/00212/2020, and FAPESP 2023/17918-2
-Abstract: For every $3$-dimensional non-Lie Leibniz algebras over complex numbers we describe the algebra of polynomial invariants and the group of automorphisms.
-As a consequence, we establish that two non-nilpotent $3$-dimensional non-Lie Leibniz algebras can be distinguished by means of the traces of degrees $\leqslant 2$ and the dimensions of groups of automorphisms.
-Keywords: Leibniz algebras, polynomial invariants, separating set, generating set.
+Polynomial invariants for $3$-dimensional Leibniz algebras  [^1] The work is supported by the FCT 2023.08031.CEECIND, UIDB/00212/2020, UIDP/00212/2020, and FAPESP 2023/17918-2
+Abstract: For each $3$-dimensional non-Lie Leibniz algebra over the complex numbers, we describe the algebra of polynomial invariants and determine its group of automorphisms.
+As a consequence, we establish that any two non-nilpotent $3$-dimensional non-Lie Leibniz algebras can be distinguished by the traces of degrees $\leqslant 2$ and by the dimensions of their automorphism groups.
+Keywords: Leibniz algebras, polynomial invariants, generating set.
 MSC2020: 13A50, 15A72, 17A32, 17A36, 20F29.
@@
 commutative and associative) algebra
-[…]
+$\mathbb{F}[\mathcal{A}^{m}]=\mathbb{F}[x_{ri}\,|\,1\leqslant r\leqslant m,\;1\leqslant i\leqslant n].$
 Fix some basis $\{e_{1},\ldots,e_{n}\}$ for $\mathcal{A}$, and for every $a\in\mathcal{A}$ denote by $(a)_{i}$ the $i-{\rm th}$ coordinate of $a$ with respect the given basis.
@@
     $\mathcal{A}=M_{n}$ is the algebra of all $n\times n$ matrices over $\mathbb{F}$, in case $\mathop{\rm char}{\mathbb{F}}=0$ or $\mathop{\rm char}{\mathbb{F}}>n$.
-Generators for the algebra of invariants $I_{m}(M_{n})$ were independently described in [36] and [32] in the case of charactristic zero, and in [11] in the case of positive characteristic.
-Minimal generating sets for $I_{m}(M_{2})$ and $I_{m}(M_{3})$ were obtained in [33, 10, 25, 26, 27] and for $I_{m}(M_{4})$ and $I_{m}(M_{5})$ with small values of $m$, in [39, 12, 9], assuming $\mathop{\rm char}\mathbb{F}=0$.
+Generators for the algebra of invariants $I_{m}(M_{n})$ were independently described in [Sibirskii68] and [Procesi76] in the case of characteristic zero, and in [Donkin92a] in the case of positive characteristic.
+Minimal generating sets for $I_{m}(M_{2})$ and $I_{m}(M_{3})$ were obtained in [Procesi84, DKZ02, Lopatin_Comm1, Lopatin_Sib, Lopatin_Comm2] and for $I_{m}(M_{4})$ and $I_{m}(M_{5})$ with small values of $m$, in [Teranishi86, Drensky_Sadikova_4x4, Djokovic07], assuming $\mathop{\rm char}\mathbb{F}=0$.
   2. $\bullet$
     $\mathcal{A}=\mathbf{O}$ is the octonion algebra, in case $\mathop{\rm char}{\mathbb{F}}\neq 2$.
-Generators for the algebra of invariants $I_{m}(\mathbf{O})$ were constructed in [35] over the field of complex numbers $\mathbb{C}$ and in [40] over an arbitrary infinite field of odd characteristic.
-In case $\mathop{\rm char}{\mathbb{F}}\neq 2$, a minimal generating set was constructed [28], using the classification of pairs of octonions from [29].
+Generators for the algebra of invariants $I_{m}(\mathbf{O})$ were constructed in [schwarz1988] over the field of complex numbers $\mathbb{C}$ and in [zubkov2018] over an arbitrary infinite field of odd characteristic.
+In case $\mathop{\rm char}{\mathbb{F}}\neq 2$, a minimal generating set was constructed [LZ_2], using the classification of pairs of octonions from [LZ_1].
   3. $\bullet$
-    $\mathcal{A}={\mathbb{A}}$ is the split Albert algebra, i.e., the exceptional simple Jordan algebra of $3\times 3$ Hermitian matrices over $\mathbf{O}$ with the symmetric multiplication $a\circ b=(ab+ba)/2$, in case $\mathop{\rm char}{\mathbb{F}}=0$ and $m\in\{1,2\}$ (see [16, 31]).
-More details about the Artin–Procesi–Iltyakov Equality can be found in [1].
+    $\mathcal{A}={\mathbb{A}}$ is the split Albert algebra, i.e., the exceptional simple Jordan algebra of $3\times 3$ Hermitian matrices over $\mathbf{O}$ with the symmetric multiplication $a\circ b=(ab+ba)/2$, in case $\mathop{\rm char}{\mathbb{F}}=0$ and $m\in\{1,2\}$ (see [Iltyakov_1995, Polikarpov_1991]).
+  4. $\bullet$
+    $\mathcal{A}$ is a two-dimensional simple algebra with a non-trivial automorphism group, in case $\mathop{\rm char}{\mathbb{F}}=0$ (see [Alvarez_Lopatin_2025]).
+More details about the Artin–Procesi–Iltyakov Equality can be found in [Alvarez_Lopatin_2025].
 1.2 Leibniz algebras
@@
 It was at first introduced by A.
-Bloh [3] and then independently by J.-L.
-Loday [23].
-Recently, they appeared in many geometric and physics applications (see, for example, [4, 22, 21, 37] and references therein).
+Bloh [Bloh_1965] and then independently by J.-L.
+Loday [L93a].
+Recently, they appeared in many geometric and physics applications (see, for example, [bonez, kotov20, KW01, strow20] and references therein).
 A systematic study of algebraic properties of Leibniz algebras was started by J.-L.
 Loday and T.
-Pirashvili [24].
+Pirashvili [L93].
 So, several classical theorems from Lie algebras theory have been extended to the Leibniz algebras case; many classification results regarding nilpotent, solvable, etc.
-Leibniz algebras have been obtained (see, for example, [13, 2, 7, 6, 14, 20, 19, 30, 38] and references therein); for results on algebraic and geometric classifications of low dimensional Leibniz algebras see [17, 18] and references therein.
+Leibniz algebras have been obtained (see, for example, [FKS_2025, AKS, deMello_Souza_2023, CK, FW, KM, kky22, MS, TX] and references therein); for results on algebraic and geometric classifications of low dimensional Leibniz algebras see [k23, KKP] and references therein.
 Given a Leibniz algebra $\mathcal{L}$, we denote by ${\rm Leib}(\mathcal{L})$ the $\mathbb{F}$-span of $a^{2}$ for all $a\in\mathcal{L}$. Since
@@
 1.3 Results
-In this section we assume that $\mathbb{F}=\mathbb{C}$ is the field of complex numbers.
+In this section, we assume that $\mathbb{F}=\mathbb{C}$ is the field of complex numbers.
 Denote by ${\bf L}_{3}$ the family of all $3$-dimensional non-Lie Leibniz algebras over the field $\mathbb{F}$.
-The classification of algebras from ${\bf L}_{3}$ modulo the action of the group of automorphisms was independently obtained by I.M.
-Rikhsiboev, I.S.
-Rakhimov [34] and by J.M.
-Casas, M.A.
-Insua, M.
-Ladra, S.
-Ladra [5].
-The equivalence of these classifications was pointed out in the Remark at the end of paper [34].
-In our paper, we use the classification from [34] (see Theorem 9).
-In Section 2, we provide key definitions of generic elements and operator traces.
-In Section 3, we present an explicit formula for calculating some operator traces (see Proposition 4) and describe the reduction of the general case to multilinear in characteristic zero (see Proposition 6).
+The classification of algebras from ${\bf L}_{3}$ modulo the action of the automorphism group was independently obtained by:
+  1. $\bullet$
+    J.M. Casas, M.A. Insua, M. Ladra, and S. Ladra [Casas_Insua_Ladra_Ladra_2012_LAA], by means of a computer program;
+  2. $\bullet$
+    I.M. Rikhsiboev and I.S. Rakhimov [Rikhsiboev_Rakhimov_2012_AIP], without the use of a computer. Unfortunately, this classification is not complete (see the proof of Theorem 9 for details).
+In our paper, we complete the classification given in [Rikhsiboev_Rakhimov_2012_AIP] in Theorem 9, and we prove its equivalence with the classification obtained in [Casas_Insua_Ladra_Ladra_2012_LAA] (see again the proof of Theorem 9).
+In Section 2, we provide the key definitions of generic elements and operator traces.
+In Section 3, we present explicit formulas for calculating certain operator traces (see Proposition 4) and describe the reduction of the general case to the multilinear case in characteristic zero (see Proposition 6).
 Assume $\mathcal{L}\in{\bf L}_{3}$.
@@
 Then in Section 5 we obtain a generating set for the algebra of invariants $I_{m}(\mathcal{L})$ (see Theorem 14).
-As a consequence, we describe when Artin–Procesi–Iltyakov Equality holds for $\mathcal{L}$ (see Corollary 16).
+As a consequence, we describe when the Artin–Procesi–Iltyakov Equality holds for $\mathcal{L}$ (see Corollary 16).
 In Corollary 18 we obtain $I_{m}(\mathcal{L})=\mathbb{F}$ for all $m\geqslant 1$ if and only if $\mathcal{L}$ is nilpotent.
@@
 Note that the problem of distinguishing low-dimensional algebras using their algebraic properties has recently been studied.
-Specifically, in [8], it was shown that any two-dimensional Jordan algebras over a finite field $\mathbb{F}$ with odd characteristic are isomorphic if and only if they satisfy the same polynomial identities.
-In [15], it was established that polynomial identities distinguish non-associative two-dimensional Novikov algebras over the complex numbers.
+Specifically, in [diniz2024isomorphism], it was shown that any two-dimensional Jordan algebras over a finite field $\mathbb{F}$ with odd characteristic are isomorphic if and only if they satisfy the same polynomial identities.
+In [PIs_Novikov_dim2_2025], it was established that polynomial identities distinguish non-associative two-dimensional Novikov algebras over the complex numbers.
 1.4 Notations
-A monomial $w=x_{r_{1},i_{1}}\cdots x_{r_{k},i_{k}}$ from $\mathbb{F}[\mathcal{A}^{m}]$ has multidegree $\mathop{\rm mdeg}(w)=(\delta_{1},\ldots,\delta_{m})\in\mathbb{N}^{m}$, where $\delta_{r}$ is the number of letters of $w$ lying in the set $\{x_{r1},\ldots,x_{rn}\}$ and =$\mathbb{N}=\{0,1,2,\ldots\}$.
+A monomial $w=x_{r_{1},i_{1}}\cdots x_{r_{k},i_{k}}$ from $\mathbb{F}[\mathcal{A}^{m}]$ has multidegree $\mathop{\rm mdeg}(w)=(\delta_{1},\ldots,\delta_{m})\in\mathbb{N}^{m}$, where $\delta_{r}$ is the number of letters of $w$ lying in the set $\{x_{r1},\ldots,x_{rn}\}$ and $\mathbb{N}=\{0,1,2,\ldots\}$.
 As an example, $\mathop{\rm mdeg}(x_{31}x_{32}x_{21}x_{12})=(1,1,2)$ for $m=3$.
@@
   2. (b)
-    It is well known that $\mathop{\rm tr}(h)$ is an invariant from $I_{m}(\mathcal{A})$ (for example, see Lemma 3.1 from [1]).
+    It is well known that $\mathop{\rm tr}(h)$ is an invariant from $I_{m}(\mathcal{A})$ (for example, see Lemma 3.1 from [Alvarez_Lopatin_2025]).
   3. (c)
@@
 $M_{\rm L}^{(i)}=(M_{ijl})_{1\leqslant l,j\leqslant n}\;\;\text{ and }\;\;M_{\rm R}^{(i)}=(M_{jil})_{1\leqslant l,j\leqslant n}.$
-Proposition 4 (Proposition 3.2 of [1]).
+Proposition 4 (Proposition 3.2 of [Alvarez_Lopatin_2025]).
 Assume that $m\geqslant 1$ and a homogeneous monomial $h\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ of degree 1 in $\chi_{0}$ satisfies the equality $h=P^{1}_{\chi_{r_{1}}}\circ\cdots\circ P^{k}_{\chi_{r_{k}}}(\chi_{0})$ for some symbols $P^{1},\ldots,P^{k}\in\{L,R\}$, $1\leqslant r_{1},\ldots,r_{k}\leqslant m$ and $k>0$.
@@
 Proposition 4 implies that for all $1\leqslant r\leqslant m$ we have
-$\mathop{\rm tr}(\chi_{r}\chi_{0})=\sum_{i,j=1}^{n}x_{ri}M_{ijj}\;\;\text{ and }\;\;\mathop{\rm tr}(\chi_{0}\chi_{r})=\sum_{i,j=1}^{n}x_{ri}M_{jij}.$ (3.1)
+[…]
 For every $1\leqslant i,i^{\prime}\leqslant n$ consider the following $n\times n$ matrices over $\mathbb{F}$:
 $M_{\rm L}^{(i,i^{\prime})}=\Big(\sum_{t=1}^{n}M_{ii^{\prime}t}M_{tjl}\Big)_{1\leqslant l,j\leqslant n}\;\;\text{ and }\;\;M_{\rm R}^{(i,i^{\prime})}=\Big(\sum_{t=1}^{n}M_{ii^{\prime}t}M_{jtl}\Big)_{1\leqslant l,j\leqslant n}.$
-Proposition 5 (Proposition 3.3 of [1]).
+Proposition 5 (Proposition 3.3 of [Alvarez_Lopatin_2025]).
 Assume that $m\geqslant 1$ and a homogeneous monomial $h\in\mathbb{F}\langle\chi_{0},\ldots,\chi_{m}\rangle$ of degree 1 in $\chi_{0}$ satisfies the equality $h=P^{1}_{h_{1}}\circ\cdots\circ P^{k}_{h_{k}}(\chi_{0})$ for some symbols $P^{1},\ldots,P^{k}\in\{L,R\}$ and monomials $h_{1},\ldots,h_{k}\in\mathbb{F}\langle\chi_{1},\ldots,\chi_{m}\rangle$ of degree 1 or 2, where $k>0$.
@@
 Given a partition ${\underline{r}}\in{\mathcal{P}}^{t}_{m}$, we define a function ${\underline{r}}|\cdot|:\{1,\ldots,t\}\to\{1,\ldots,m\}$ by
-${\underline{r}}|l|=j\text{ if and only if }r_{1}+\cdots+r_{j-1}+1\leqslant l\leqslant r_{1}+\cdots+r_{j},$
+${\underline{r}}|l|=j\text{ if and only if }\sum_{1\leqslant i\leqslant j-1}r_{i}+1\leqslant l\leqslant\sum_{1\leqslant i\leqslant j}r_{i},$
 for all $1\leqslant l\leqslant t$. For a partition ${\underline{r}}\in{\mathcal{P}}^{t}_{m}$ we also define a homomorphism of $\mathbb{F}$-algebras as follows:
@@
 The following proposition and lemma are well known and can easily be proven.
-Proposition 6.
+Proposition 6 (cf. II.5, Theorem 2.5A of [Weyl_book]).
 The algebra of invariants $\mathbb{F}[\mathcal{A}^{m}]^{G}$ is generated by
 $\{\pi_{{\underline{r}}}(f)\,|\,f\in\mathbb{F}[\mathcal{A}^{t}]^{G}\text{ is multilinear},\;{\underline{r}}\in{\mathcal{P}}^{t}_{m},\;t\geqslant m\}.$
-Lemma 7.
+Lemma 7 (cf. Theorem 5.14 of [Bruns_Gubeladze_book_2009]).
 If the group $G$ is diagonal, i.e., all elements of $G\leqslant{\rm GL}_{n}$ are diagonal matrices, then the algebra of invariants $\mathbb{F}[\mathcal{A}^{m}]^{G}$ is generated by some monomials from $\mathbb{F}[\mathcal{A}^{m}]$.
-The following lemma follows immediately from a partial case of Theorem 7.6 from [1] for the algebra ${\mathbf{A}}_{2}$.
-Lemma 8.
+The following lemma follows immediately from a partial case of Theorem 7.6 from [Alvarez_Lopatin_2025] for the algebra ${\mathbf{A}}_{2}$.
+Lemma 8 (see Theorem 7.6 of [Alvarez_Lopatin_2025]).
 Assume that $\mathcal{V}$ is a two-dimensional vector space with a basis $\{e_{1},e_{2}\}$ and let $G\leqslant{\rm GL}_{2}$ be the group of all matrices
 $\left(\begin{array}[]{cc}1&\alpha\\ 0&1\\ \end{array}\right),$
-[…]
+where $\alpha\in\mathbb{F}$.
+As in Section 1.1, the coordinate ring $\mathbb{F}[\mathcal{V}^{m}]=\mathbb{F}[x_{r1},x_{r2}\,|\,1\leqslant r\leqslant m]$ of $\mathcal{V}^{m}$ is a $G$-module.
+Then the algebra of invariants $\mathbb{F}[\mathcal{V}^{m}]^{G}$ is generated by
 $1,x_{12},\ldots,x_{m2},\;x_{r1}x_{s2}-x_{r2}x_{s1}\;\;(1\leqslant r<s\leqslant m).$
@@
 We use notations for the tableau of multiplication of an algebra from Section 3.1.
-Theorem 9 (Rikhsiboev, Rakhimov [34], Theorem 2).
+Theorem 9.
 Each algebra $\mathcal{L}\in{\bf L}_{3}$ is isomorphic to one and only one algebra from the following list, where algebras are given by their tableaux of multiplication with respect to some basis $\{e_{1},e_{2},e_{3}\}$:
-$\mathcal{L}_{1}$ :::: $\left(\begin{array}[]{ccc}0&0&-2e_{1}\\ 0&e_{1}&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{2}^{\lambda}$ :::: $\left(\begin{array}[]{ccc}0&0&\lambda e_{1}\\ 0&0&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{3}$ :::: $\left(\begin{array}[]{ccc}0&0&0\\ 0&0&-e_{2}\\ 0&e_{2}&e_{1}\\ \end{array}\right),$
-$\mathcal{L}_{4}^{\lambda}$ :::: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&e_{1}\\ 0&0&\lambda e_{1}\\ \end{array}\right),$ $\mathcal{L}_{5}$ :::: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{6}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&e_{1}\\ 0&0&0\\ \end{array}\right),$
-$\mathcal{L}_{7}^{\lambda}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&\lambda e_{1}+e_{2}\\ 0&0&0\\ \end{array}\right),$ $\mathcal{L}_{8}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{9}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{1}+e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$
+$\mathcal{L}_{1}$ :: $\left(\begin{array}[]{ccc}0&0&-2e_{1}\\ 0&e_{1}&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{2}^{{\lambda\neq 0}}$ :: $\left(\begin{array}[]{ccc}0&0&\lambda e_{1}\\ 0&0&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{3}$ :: $\left(\begin{array}[]{ccc}0&0&0\\ 0&0&-e_{2}\\ 0&e_{2}&e_{1}\\ \end{array}\right),$
+$\mathcal{L}_{4}^{\lambda}$ :: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&e_{1}\\ 0&0&\lambda e_{1}\\ \end{array}\right),$ $\mathcal{L}_{5}$ :: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{6}$ :: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&e_{1}\\ 0&0&0\\ \end{array}\right),$
+$\mathcal{L}_{7}^{\lambda}$ :: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&\lambda e_{1}+e_{2}\\ 0&0&0\\ \end{array}\right),$ $\mathcal{L}_{8}$ :: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{9}$ :: $\left(\begin{array}[]{ccc}0&0&e_{1}+e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$
+$\mathcal{L}_{10}$ :: $\left(\begin{array}[]{ccc}0&0&e_{1}\\ 0&0&e_{2}\\ 0&0&0\\ \end{array}\right),$ $\mathcal{L}_{11}$ :: $\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right).$
 where $\lambda\in\mathbb{F}$.
+Proof.
+A classification of 3-dimensional non-Lie Leibniz algebras over complex numbers was obtained by Rikhsiboev, Rakhimov in Theorem 2 of [Rikhsiboev_Rakhimov_2012_AIP], where it was claimed that any algebra from ${\bf L}_{3}$ belongs to the following list: $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{4}^{\lambda}$, $\mathcal{L}_{5}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$, $\mathcal{L}_{8}$, $\mathcal{L}_{9}$ with $\lambda\in\mathbb{F}$.
+Unfortunately, this classification is not complete:
+  1. $\bullet$
+    The algebra $\mathcal{L}_{2}^{0}$ is a Lie algebra, therefore, we should assume that $\lambda$ is non-zero in case $\mathcal{L}_{2}^{\lambda}$.
+  2. $\bullet$
+    On page 361 of the proof of Theorem 2 of [Rikhsiboev_Rakhimov_2012_AIP] it was claimed that $\mathcal{L}_{10}$ is isomorphic to $\mathcal{L}_{6}$, which is not the case.
+Actually, as it follows from Corollary 19 (see below), $\mathcal{L}_{10}$ is not isomorphic to any other algebra from the formulation of Theorem 9.
+  3. $\bullet$
+    The algebra $\mathcal{L}_{11}$ was missed in Case 2.1 of the proof of Theorem 2 of [Rikhsiboev_Rakhimov_2012_AIP].
+To show that $\mathcal{L}_{11}$ is not isomorphic to any other algebra from the formulation of Theorem 9, we note that the only commutative algebras from the formulation of Theorem 9 are $\mathcal{L}_{5}$ and $\mathcal{L}_{11}$, but $\dim{\rm Ann}^{\rm R}(\mathcal{L}_{5})=1$ and $\dim{\rm Ann}^{\rm R}(\mathcal{L}_{11})=2$, where the right annihilator of an algebra $\mathcal{A}$ is Ann^R(A) ={a∈A —  ba=0 for all b∈A}.
+To complete the proof, we will prove that the classification from Theorem 9 is equivalent to the computer-based classification of 3-dimensional non-Lie Leibniz algebras over complex numbers by J.M.
+Casas, M.A.
+Insua, M.
+Ladra, S.
+Ladra [Casas_Insua_Ladra_Ladra_2012_LAA].
+For $\delta\in\mathbb{F}$, we denote by
+$\mathcal{L}_{2a}^{\delta},\;\mathcal{L}_{2b},\;\mathcal{L}_{2c},\;\mathcal{L}_{2d},\;\mathcal{L}_{2e}^{\delta}\;(\delta\neq 0),\;\mathcal{L}_{2f},\;\mathcal{L}_{2g},\;\mathcal{L}_{3a}^{\delta}\;(\delta\in\widetilde{\mathbb{F}}),\;\mathcal{L}_{3b},\;\mathcal{L}_{3c},\;\mathcal{L}_{3d}$ (4.1)
+the 3-dimensional Leibniz algebras from page 3752 of [Casas_Insua_Ladra_Ladra_2012_LAA], where
+$\widetilde{\mathbb{F}}=\{\alpha\in\mathbb{C}\,|\,|\alpha|>1\}\cup\{\alpha\in\mathbb{C}\,|\,|\alpha|=1,\;0\leqslant{\rm arg}(\alpha)\leqslant\pi\}.$
+[…]
+  1. $\bullet$
+    for each line, the algebra from the first column is isomorphic to the algebra from the second column;
+  2. $\bullet$
+    the tableau of multiplication of the algebra from the first column considered with respect to the ordered basis from the third column is equal to the tableau of multiplication of the algebra from the second column.
+$\mathcal{L}_{1}$ $\mathcal{L}_{2g}$ $\{e_{1},e_{1}+e_{2},-e_{2}-e_{3}\}$
+$\mathcal{L}_{2}^{\lambda\neq 0}$ $\mathcal{L}_{2e}^{-\lambda}$ $\{e_{1},e_{2},-e_{3}\}$
+$\mathcal{L}_{3}$ $\mathcal{L}_{2f}$ $\{e_{1},e_{2},-e_{3}\}$
+$\mathcal{L}_{4}^{\lambda\neq 0}$ $\mathcal{L}_{2a}^{\lambda}$ $\{e_{1},e_{3},e_{2}\}$
+$\mathcal{L}_{4}^{0}$ $\mathcal{L}_{2a}^{0}$ $\{e_{1},e_{3},e_{2}\}$
+$\mathcal{L}_{5}$ $\mathcal{L}_{2c}$ $\{e_{1},e_{2},e_{3}\}$
+$\mathcal{L}_{6}$ $\mathcal{L}_{3a}^{-1}$ $\{e_{1}+e_{2},e_{2}-e_{1},-e_{3}\}$
+$\mathcal{L}_{7}^{\lambda\neq-\frac{1}{4},0}$ $\mathcal{L}_{3a}^{\lambda^{\prime}}$ $\{\xi_{1}e_{1}+e_{2},\xi_{2}e_{1}+e_{2},\nu e_{3}\}$
+$\mathcal{L}_{7}^{-\frac{1}{4}}$ $\mathcal{L}_{3b}$ $\{e_{2},-\frac{e_{1}}{2}+e_{2},2e_{3}\}$
+$\mathcal{L}_{7}^{0}$ $\mathcal{L}_{2d}$ $\{e_{2},e_{1}-e_{2},e_{3}\}$
+$\mathcal{L}_{8}$ $\mathcal{L}_{3c}$ $\{e_{1},e_{2},e_{3}\}$
+$\mathcal{L}_{9}$ $\mathcal{L}_{3d}$ $\{2e_{1}+e_{2},2e_{1}+2e_{2},e_{1}+e_{3}\}$
+$\mathcal{L}_{10}$ $\mathcal{L}_{3a}^{1}$ $\{e_{1},e_{2},e_{3}\}$
+$\mathcal{L}_{11}$ $\mathcal{L}_{2b}$ $\{e_{1},e_{2},e_{3}\}$
+Here,
+  1. $\bullet$
+[…]
+  2. $\bullet$
+    $\nu=\frac{2\xi_{1}+1}{\xi_{1}-2\lambda}$.
+[…]
+$\xi_{1}\xi_{2}=-\lambda\;\text{ and }\;\xi_{1}+\xi_{2}=-1.$ (4.2)
+Since $\lambda\neq-1/4$, we have
+$\xi_{1}\neq 2\lambda\;\text{ and }\;\xi_{2}\neq 2\lambda.$ (4.3)
+[…]
+Formulas (4.3) imply that $\nu$ is well-defined and $\nu\neq 0$. Hence,
+$v_{1}=\xi_{1}e_{1}+e_{2},\;\;v_{2}=\xi_{2}e_{1}+e_{2},\;\;v_{3}=\nu e_{3}$
+is a basis for $\mathcal{L}$.
+Compute the tableau of multiplication for $\mathcal{L}$ with respect to the ordered basis $\{v_{1},v_{2},v_{3}\}$:
+  1. $\bullet$
+    Using formulas (4.2) it is not difficult to see that $\nu\lambda=\lambda^{\prime}\xi_{1}$ and $\nu(\xi_{1}+1)=\lambda^{\prime}$.
+Hence, v_1v_3=ξ_1νe_2 + ν(λe_1+e_2) = νλe_1 + ν(ξ_1+1) e_2 = λξ_1 e_1 + λ’ e_2 = λ’ v_1.
+  2. $\bullet$
+    Using formulas (4.2) it is not difficult to see that $\nu\lambda=\xi_{2}$ and $\nu(\xi_{2}+1)=1$. Hence, v_2v_3=ξ_2νe_2 + ν(λe_1+e_2) = νλe_1 + ν(ξ_2+1) e_2 = ξ_2 e_1 + e_2 = v_2.
+  3. $\bullet$
+    others products $v_{i}v_{j}$ are zero.
+Therefore, the tableau of multiplication for $\mathcal{L}$ with respect to the ordered basis $\{v_{1},v_{2},v_{3}\}$ is
+$\left(\begin{array}[]{ccc}0&0&\lambda^{\prime}v_{1}\\ 0&0&v_{2}\\ 0&0&0\\ \end{array}\right).$
+Hence, $\mathcal{L}\simeq\mathcal{L}_{3a}^{\lambda^{\prime}}$. The following claim, which can be proven by straightforward computations, completes the proof:
+[…]
+where $\xi_{1},\xi_{2}\in\mathbb{F}$ are pairwise different roots of the equation $x^{2}+x-\lambda=0$.
 Theorem 10.
 The groups of automorphisms of algebras from ${\bf L}_{3}$ are the following set of matrices:
-$\mathcal{L}_{1}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{5}^{2}&\alpha_{5}\alpha_{6}&\alpha_{6}^{2}/2\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
-$\mathcal{L}_{2}^{\lambda\neq 0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&0\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1},\alpha_{5}\neq 0;$
-$\mathcal{L}_{2}^{0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1},\alpha_{5}\neq 0;$
-$\mathcal{L}_{3}$ :::: $g=\left(\begin{array}[]{ccc}1&0&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
-$\mathcal{L}_{4}^{\lambda\neq 0}$ :::: $g_{1}=\left(\begin{array}[]{ccc}\frac{\alpha_{6}^{2}+\alpha_{6}\alpha_{9}+\lambda\alpha_{9}^{2}}{\lambda}&\alpha_{2}&\alpha_{3}\\ 0&\frac{\alpha_{6}}{\lambda}+\alpha_{9}&\alpha_{6}\\ 0&-\frac{\alpha_{6}}{\lambda}&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{6}\neq 0,\;\det(g_{1})\neq 0,$
+$\mathcal{L}_{1}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{5}^{2}&\alpha_{5}\alpha_{6}&\alpha_{6}^{2}/2\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
+$\mathcal{L}_{2}^{{\lambda\neq 0}}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&0\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1},\alpha_{5}\neq 0;$
+$\mathcal{L}_{3}$ :: $g=\left(\begin{array}[]{ccc}1&0&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
+$\mathcal{L}_{4}^{\lambda\neq 0}$ :: $g_{1}=\left(\begin{array}[]{ccc}\frac{\alpha_{6}^{2}+\alpha_{6}\alpha_{9}+\lambda\alpha_{9}^{2}}{\lambda}&\alpha_{2}&\alpha_{3}\\ 0&\frac{\alpha_{6}}{\lambda}+\alpha_{9}&\alpha_{6}\\ 0&-\frac{\alpha_{6}}{\lambda}&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{6}\neq 0,\;\det(g_{1})\neq 0,$
 $g_{2}=\left(\begin{array}[]{ccc}\alpha_{5}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&0\\ 0&0&\alpha_{5}\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
-$\mathcal{L}_{4}^{0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{5}(\alpha_{5}+\alpha_{8})&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&0\\ 0&\alpha_{8}&\alpha_{5}+\alpha_{8}\\ \end{array}\right)\text{ for }\alpha_{5}+\alpha_{8}\neq 0;$
-$\mathcal{L}_{5}$ :::: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{5}^{2}+\alpha_{6}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&-\alpha_{6}&\alpha_{5}\\ \end{array}\right)\text{ for }a_{5}\neq 0,\;\alpha_{5}^{2}+\alpha_{6}^{2}\neq 0,$
+$\mathcal{L}_{4}^{0}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{5}(\alpha_{5}+\alpha_{8})&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&0\\ 0&\alpha_{8}&\alpha_{5}+\alpha_{8}\\ \end{array}\right)\text{ for }\alpha_{5}+\alpha_{8}\neq 0;$
+$\mathcal{L}_{5}$ :: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{5}^{2}+\alpha_{6}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&-\alpha_{6}&\alpha_{5}\\ \end{array}\right)\text{ for }a_{5}\neq 0,\;\alpha_{5}^{2}+\alpha_{6}^{2}\neq 0,$
 $g_{2}=\left(\begin{array}[]{ccc}\alpha_{5}^{2}+\alpha_{6}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&\alpha_{6}&-\alpha_{5}\\ \end{array}\right)\text{ for }a_{5}\neq 0,\;\alpha_{5}^{2}+\alpha_{6}^{2}\neq 0,$
 $g_{3}=\left(\begin{array}[]{ccc}\alpha_{6}^{2}&\alpha_{2}&\alpha_{3}\\ 0&0&\alpha_{6}\\ 0&\pm\alpha_{6}&0\\ \end{array}\right)\text{ for }\alpha_{6}\neq 0;$
-$\mathcal{L}_{6}$ :::: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{1}&\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}\neq\alpha_{4}^{2};$
+$\mathcal{L}_{6}$ :: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{1}&\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}\neq\alpha_{4}^{2};$
 $g_{2}=\left(\begin{array}[]{ccc}\alpha_{1}&-\alpha_{4}&0\\ \alpha_{4}&-\alpha_{1}&0\\ 0&0&-1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}\neq\alpha_{4}^{2};$
-$\mathcal{L}_{7}^{\lambda\neq 0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&\lambda\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}+\alpha_{1}\alpha_{4}-\lambda\alpha_{4}^{2}\neq 0;$
-$\mathcal{L}_{7}^{0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&\alpha_{3}\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&-\alpha_{3}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}(\alpha_{1}+\alpha_{4})\neq 0;$
-$\mathcal{L}_{8}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{9}^{2}&0&\alpha_{3}\\ \alpha_{3}\alpha_{9}&\alpha_{9}^{3}&\alpha_{6}\\ 0&0&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{9}\neq 0;$
-$\mathcal{L}_{9}$ :::: $g=\left(\begin{array}[]{ccc}1+\alpha_{3}&0&\alpha_{3}\\ \alpha_{3}&1&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{3}\neq-1.$
+$\mathcal{L}_{7}^{\lambda\neq 0}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{1}&\lambda\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}+\alpha_{1}\alpha_{4}-\lambda\alpha_{4}^{2}\neq 0;$
+$\mathcal{L}_{7}^{0}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&\alpha_{3}\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&-\alpha_{3}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}(\alpha_{1}+\alpha_{4})\neq 0;$
+$\mathcal{L}_{8}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{9}^{2}&0&\alpha_{3}\\ \alpha_{3}\alpha_{9}&\alpha_{9}^{3}&\alpha_{6}\\ 0&0&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{9}\neq 0;$
+$\mathcal{L}_{9}$ :: $g=\left(\begin{array}[]{ccc}1+\alpha_{3}&0&\alpha_{3}\\ \alpha_{3}&1&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{3}\neq-1;$
+$\mathcal{L}_{10}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{1}&\alpha_{2}&0\\ \alpha_{4}&\alpha_{5}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}\alpha_{5}-\alpha_{2}\alpha_{4}\neq 0;$
+$\mathcal{L}_{11}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{9}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{5}\alpha_{9}\neq 0.$
 Here $\alpha_{1},\ldots,\alpha_{9}\in\mathbb{F}$.
@@
   2. 2.
-    Assume $\mathcal{L}=\mathcal{L}_{2}^{\lambda}$ for some $\lambda\in\mathbb{F}$.
-    1. (a)
-      Let $\lambda\neq 0$.
+    Assume $\mathcal{L}=\mathcal{L}_{2}^{\lambda}$ for some non-zero $\lambda\in\mathbb{F}$.
 Consequently considering $g_{13}=0$, $g_{23}=0$, $g_{32}=0$ we obtain $\alpha_{7}=0$, $\alpha_{8}=0$, $\alpha_{2}=0$, respectively.
 Since $\det(g)=\alpha_{1}\alpha_{5}\alpha_{9}\neq 0$, we consequently consider $g_{31}=0$, $g_{33}=0$, $g_{13}=0$ to obtain $\alpha_{4}=0$, $\alpha_{3}=0$, $\alpha_{9}=1$, respectively.
-    2. (b)
-      Let $\lambda=0$.
-Since $g_{32}=0$, we have $\alpha_{2}=0$ and $\alpha_{8}=0$.
-Since $\det(g)=-\alpha_{3}\alpha_{5}\alpha_{7}+\alpha_{1}\alpha_{5}\alpha_{9}$, we have $\alpha_{5}\neq 0$.
-Since $g_{12}=0$, we have $\alpha_{7}=0$.
-Thus $\det(g)=\alpha_{1}\alpha_{5}\alpha_{9}\neq 0$ and we consequently consider $g_{13}=0$, $g_{23}=0$ to obtain $\alpha_{4}=0$, $\alpha_{9}=1$, respectively.
-The required statement is proven.
   3. 3.
@@
     In case $\alpha_{9}=-\alpha_{3}$ we have $\det(g)=0$; a contradiction. Therefore, $\alpha_{9}\neq-\alpha_{3}$ and $g_{13}=0$ implies that $\alpha_{9}=1$. The required statement is proven.
+  10. 10.
+    Assume $\mathcal{L}=\mathcal{L}_{10}$.
+Equality $g_{13}=0$ implies that $\alpha_{7}=0$.
+Applying equality $g_{23}=0$, we can see that $\alpha_{8}=0$.
+Note that $\alpha_{9}\neq 0$, since $\det(g)\neq 0$.
+Hence, it follows from $g_{33}=0$ that $\alpha_{3}=\alpha_{6}=0$.
+    1. (a)
+      Let $\alpha_{9}\neq 1$. Then $g_{13}=0$ implies $\alpha_{1}=\alpha_{4}=0$; hence $\det(g)=0$, a contradiction.
+    2. (b)
+      Let $\alpha_{9}=1$. Then the required statement holds.
+  11. 11.
+    Assume $\mathcal{L}=\mathcal{L}_{11}$.
+Equality $g_{11}=0$ implies that $\alpha_{7}=0$.
+Applying equality $g_{22}=0$, we can see that $\alpha_{8}=0$.
+Hence, it follows from equality $g_{33}=0$ that $\alpha_{4}=0$ and $\alpha_{1}=\alpha_{9}^{2}$.
+The required statement is proven.
 Corollary 11.
-Given $\mathcal{L}\in{\bf L}_{3}$, the group $\mathop{\rm Aut}(\mathcal{L})$ has dimension $2$, $3$ or $4$ as an affine variety, where each of the possibilities holds for some algebra $\mathcal{L}\in{\bf L}_{3}$.
+Given $\mathcal{L}\in{\bf L}_{3}$, the group $\mathop{\rm Aut}(\mathcal{L})$ has dimension $2\leqslant d\leqslant 5$ as an affine variety, where each of the possibilities holds for some algebra $\mathcal{L}\in{\bf L}_{3}$.
+Proof.
+For each $\mathcal{L}\in{\bf L}_{3}$, the group $\mathop{\rm Aut}(\mathcal{L})$ is the following affine variety:
+$\left\{\Big(b_{1},\ldots,b_{d},f_{1}({\underline{b}}),\ldots,f_{r}({\underline{b}})\Big)\,\Big|\,b_{1},\ldots,b_{d}\in\mathbb{F}\text{ and }h_{1}({\underline{b}})\neq 0,\ldots,h_{s}({\underline{b}})\neq 0\right\}$
+for some polynomials $f_{1},\ldots,f_{r},h_{1},\ldots,h_{s}$ from $\mathbb{F}[x_{1},\ldots,x_{d}]$ with $r,s\geqslant 0$, where $\underline{b}$} stands for $(b_{1},\ldots,b_{d})$.
+Then the dimension of $\mathop{\rm Aut}(\mathcal{L})$ is $d$.
+Applying this reasoning together with Theorem 10, we complete the proof.
+∎
 An automorphism $\delta$ of an algebra $\mathcal{A}$ is called diagonal if there exists some basis $\{e_{1},\ldots,e_{n}\}$ of $\mathcal{A}$ such that $\varphi(e_{i})=\alpha_{i}e_{i}$ for some $\alpha_{i}\in\mathbb{F}$.
@@
 Remark 12.
-[…]
+If there are $a_{1},\ldots,a_{n}\in\mathbb{N}$ such that $\delta={\rm diag}\{\alpha^{a_{1}},\ldots,\alpha^{a_{n}}\}$ is an automorphism of $\mathcal{A}$ for all $\alpha\in\mathbb{F}$ and $a_{i_{0}}>0$ for some $1\leqslant i_{0}\leqslant n$, then the algebra of invariants $I_{m}(\mathcal{L})$ lies in $\mathbb{F}[x_{ri}\,|\,1\leqslant r\leqslant m,\;1\leqslant i\leqslant n,\;i\neq i_{0}]$.
+As an example, if ${\rm diag}\{\alpha,1,\ldots,1\}$ is an automorphism of $\mathcal{A}$ for all $\alpha\in\mathbb{F}$, then $I_{m}(\mathcal{L})$ lies in $\mathbb{F}[x_{ri}\,|\,2\leqslant r\leqslant m,\;1\leqslant i\leqslant n]$.
 Theorem 10 implies the following lemma.
@@
 Here is a list of some diagonal automorphisms $\delta$:
-$\mathcal{L}_{1}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha,1\},$ $\mathcal{L}_{2}^{\lambda}$ :::: $\delta={\rm diag}\{\alpha,\beta,1\},$
-$\mathcal{L}_{3}$ :::: $\delta={\rm diag}\{1,\alpha,1\},$ $\mathcal{L}_{4}^{\lambda}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha,\alpha\},$
-$\mathcal{L}_{5}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha,\pm\alpha\},$ $\mathcal{L}_{6}$ :::: $\delta={\rm diag}\{\alpha,\alpha,1\},$
-$\mathcal{L}_{7}^{\lambda}$ :::: $\delta={\rm diag}\{\alpha,\alpha,1\},$ $\mathcal{L}_{8}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha^{3},\alpha\},$
+$\mathcal{L}_{1}$ :: $\delta={\rm diag}\{\alpha^{2},\alpha,1\},$ $\mathcal{L}_{2}^{{\lambda\neq 0}}$ :: $\delta={\rm diag}\{\alpha,\beta,1\},$
+$\mathcal{L}_{3}$ :: $\delta={\rm diag}\{1,\alpha,1\},$ $\mathcal{L}_{4}^{\lambda}$ :: $\delta={\rm diag}\{\alpha^{2},\alpha,\alpha\},$
+$\mathcal{L}_{5}$ :: $\delta={\rm diag}\{\alpha^{2},\alpha,\pm\alpha\},$ $\mathcal{L}_{6}$ :: $\delta={\rm diag}\{\alpha,\alpha,1\},$
+$\mathcal{L}_{7}^{\lambda}$ :: $\delta={\rm diag}\{\alpha,\alpha,1\},$ $\mathcal{L}_{8}$ :: $\delta={\rm diag}\{\alpha^{2},\alpha^{3},\alpha\},$
+$\mathcal{L}_{10}$ :: $\delta={\rm diag}\{\alpha,\beta,1\},$ $\mathcal{L}_{11}$ :: $\delta={\rm diag}\{\alpha^{2},\beta,\alpha\},$
 for all $\lambda\in\mathbb{F}$ and all non-zero $\alpha,\beta\in\mathbb{F}$.
@@
   1. $\bullet$
-    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{\lambda})$, $I_{m}(\mathcal{L}_{7})$ are all equal to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
+    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{{\lambda\neq 0}})$, $I_{m}(\mathcal{L}_{7})$, $I_{m}(\mathcal{L}_{10})$ are all equal to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
   2. $\bullet$
@@
   3. $\bullet$
-    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})=\mathbb{F}$,
+    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})={I_{m}(\mathcal{L}_{11})}=\mathbb{F}$,
   4. $\bullet$
@@
   1. $\bullet$
-    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{\lambda})$, $I_{m}(\mathcal{L}_{6})$, $I_{m}(\mathcal{L}_{7})$ belong to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
+    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{{\lambda\neq 0}})$, $I_{m}(\mathcal{L}_{6})$, $I_{m}(\mathcal{L}_{7})$, $I_{m}(\mathcal{L}_{10})$ belong to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
   2. $\bullet$
@@
   3. $\bullet$
-    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})=\mathbb{F}$.
-Theorem 10 implies that $z_{1},\ldots,z_{m}\in I_{m}(\mathcal{L})$ for every $\mathcal{L}$ from the following list: $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{7}$, $\mathcal{L}_{9}$.
-Thus, the statement of the theorem holds for algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{4}^{\lambda}$, $\mathcal{L}_{5}$, $\mathcal{L}_{7}$, $\mathcal{L}_{8}$.
+    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})={I_{m}(\mathcal{L}_{11})}=\mathbb{F}$.
+Theorem 10 implies that $z_{1},\ldots,z_{m}\in I_{m}(\mathcal{L})$ for every $\mathcal{L}$ from the following list: $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{{\lambda\neq 0}}$, $\mathcal{L}_{3}$, $\mathcal{L}_{7}$, $\mathcal{L}_{9}$, $\mathcal{L}_{10}$.
+Thus, the statement of the theorem holds for algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{{\lambda\neq 0}}$, $\mathcal{L}_{4}^{\lambda}$, $\mathcal{L}_{5}$, $\mathcal{L}_{7}$, $\mathcal{L}_{8}$, $\mathcal{L}_{10}$, $\mathcal{L}_{11}$.
 Consider the algebra $\mathcal{L}_{3}$.
@@
 Restricting the group $\mathop{\rm Aut}(\mathcal{L}_{6})$ to $\mathbb{F}\text{-span}\{e_{3}\}$ and applying Lemma 7 we conclude the proof about $\mathcal{L}_{6}$.
-Consider the algebra $\mathcal{L}_{9}$.
-Using formlas (5.1) we can see that $(x_{r}-y_{r})z_{s}-z_{r}(x_{s}-y_{s})$ lies in $I_{m}(\mathcal{L}_{9})$.
-For every $\alpha\in\mathbb{F}\backslash\{-1\}$ consider an automorphism
+[…]
 $g=\left(\begin{array}[]{ccc}1+\alpha&0&\alpha\\ \alpha&1&0\\ 0&0&1\\ \end{array}\right)$
@@
 $ge_{1}^{\prime}=e_{1}^{\prime}+\alpha e_{3},\;\;ge_{2}=\alpha e^{\prime}_{1}+(\alpha+1)e_{2},\;\;ge_{3}=e_{3}.$ (5.2)
-[…]
+For $x_{r}^{\prime}=x_{r}-y_{r}$ we have that $\mathbb{F}[\mathcal{L}^{m}]=\mathbb{F}[x^{\prime}_{r},y_{r},z_{r}\,|\,1\leqslant r\leqslant m]$, and until the end of this proof we consider elements of $\mathbb{F}[\mathcal{L}^{m}]$ as linear combinations of monomials in $x^{\prime}_{r}$, $y_{r}$, $z_{r}$.
 For each $m\geqslant 0$, denote by $\Omega_{m}$ the set of all monomials $w=w_{1}\cdots w_{m}\in\mathbb{F}[\mathcal{L}^{m}]$ of multidegree $1^{m}$ for some $w_{r}\in\{x^{\prime}_{r},y_{r},z_{r}\}$, where $1\leqslant r\leqslant m$.
 Given $w\in\Omega_{m}$, denote
-$\begin{array}[]{rcl}\deg_{x^{\prime}}(w)&=&\deg_{x^{\prime}_{1}}(w)+\cdots+\deg_{x_{m}}(w),\\ \deg_{y}(w)&=&\deg_{y_{1}}(w)+\cdots+\deg_{y_{m}}(w),\\ \deg_{z}(w)&=&\deg_{z_{1}}(w)+\cdots+\deg_{z_{m}}(w).\\ \end{array}$
+[…]
 As an example, for $m=3$ we have $\deg_{y}(z_{1}x^{\prime}_{2}y_{3})=1$ and $\deg_{y}(x^{\prime}_{1}y_{2}y_{3})=2$.
@@
 As an example, for $m=3$ we have
-$x_{1}^{\prime}x_{2}^{\prime}y_{3}<z_{1}y_{2}y_{3},\;\;\;\;x_{1}^{\prime}x_{2}^{\prime}y_{3}<z_{1}x^{\prime}_{2}y_{3},\;\;\text{ and }\;\;y_{1}z_{2}x^{\prime}_{3}<y_{1}x^{\prime}_{2}y_{3}.$
+[…]
 For a multilinear $f=\sum_{j}\beta_{j}w_{j}\in\mathbb{F}[\mathcal{L}^{m}]$ with non-zero $\beta_{j}\in\mathbb{F}$ and $w_{j}\in\Omega_{m}$ we say that $w_{j_{0}}$ is the highest term for $f$, if there is no $j\neq j_{0}$ with $w_{j}>w_{j_{0}}$.
@@
 $\left(\begin{array}[]{cc}1&\alpha\\ 0&1\\ \end{array}\right)$
-[…]
+for all $\alpha\in\mathbb{F}$.
+Since the coordinate ring of $V$ is $\mathbb{F}[x^{\prime}_{r},z_{r}\,|\,1\leqslant r\leqslant m]$, applying Lemma 8 we conclude the proof of the statement about $\mathcal{L}_{9}$.
 6 Trace invariants
@@
 $\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=-z_{r}z_{s},$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=5z_{r}z_{s}.$
-$\mathcal{L}_{2}^{\lambda}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=z_{r},$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=(\lambda-1)z_{r},$
+$\mathcal{L}_{2}^{{\lambda\neq 0}}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=z_{r},$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=(\lambda-1)z_{r},$
 $\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=z_{r}z_{s},$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=-z_{r}z_{s},$
@@
 $\mathop{\rm tr}((\chi_{r}\chi_{s})\chi_{0}))=0,$ $\mathop{\rm tr}(\chi_{0}(\chi_{r}\chi_{s}))=0.$
+$\mathcal{L}_{10}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=2z_{r},$
+$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=0,$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=0,$
+$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=0,$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=2z_{r}z_{s}.$
+$\mathcal{L}_{11}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=0,$
+$\mathop{\rm tr}(\chi_{r}(\chi_{s}\chi_{0}))=0,$ $\mathop{\rm tr}((\chi_{s}\chi_{0})\chi_{r})=0,$
+$\mathop{\rm tr}(\chi_{r}(\chi_{0}\chi_{s}))=0,$ $\mathop{\rm tr}((\chi_{0}\chi_{s})\chi_{r})=0.$
 Proof.
@@
   1. $\bullet$
-    ${\rm ncl}(\mathcal{L}_{4}^{\lambda})={\rm ncl}(\mathcal{L}_{5})=3$ for all $\lambda\in\mathbb{F}$;
+    ${\rm ncl}(\mathcal{L}_{4}^{\lambda})={\rm ncl}(\mathcal{L}_{5})={{\rm ncl}(\mathcal{L}_{11})}=3$ for all $\lambda\in\mathbb{F}$;
   2. $\bullet$
@@
   3. $\bullet$
-    the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$, $\mathcal{L}_{9}$ are not nilpotent for all $\lambda\in\mathbb{F}$.
+    the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{{\lambda\neq 0}}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$, $\mathcal{L}_{9}$, $\mathcal{L}_{10}$ are not nilpotent for all $\lambda\in\mathbb{F}$.
 Proof.
-For the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$ the elements $(\cdots((e_{2}e_{3})e_{3})\cdots)e_{3}$ are non-zero.
-Similarly, $(\cdots((e_{1}e_{3})e_{3})\cdots)e_{3}\neq 0$ in $\mathcal{L}_{9}$.
+For the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{{\lambda\neq 0}}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$ the elements $(\cdots((e_{2}e_{3})e_{3})\cdots)e_{3}$ are non-zero.
+Similarly, $(\cdots((e_{1}e_{3})e_{3})\cdots)e_{3}\neq 0$ in $\mathcal{L}_{9}$ and in $\mathcal{L}_{10}$.
 Since $(e_{3}e_{3})e_{3}=e_{2}\neq 0$ in $\mathcal{L}_{8}$, we can see that ${\rm ncl}(\mathcal{L}_{8})=4$.
-The cases of $\mathcal{L}_{4}^{\lambda}$ and $\mathcal{L}_{5}$ are trivial.
+The cases of $\mathcal{L}_{4}^{\lambda}$, $\mathcal{L}_{5}$ and $\mathcal{L}_{11}$ are trivial.
 ∎
@@
 Then
-$\mathcal{L}\simeq\mathcal{L}_{1}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=2.$
-$\mathcal{L}\simeq\mathcal{L}_{2}^{\lambda\neq 0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=(\lambda-1)\mathop{\rm tr}(\chi_{1}\chi_{0}).$
-$\mathcal{L}\simeq\mathcal{L}_{2}^{0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=4.$
-$\mathcal{L}\simeq\mathcal{L}_{3}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=-\mathop{\rm tr}(\chi_{1}\chi_{0}).$
-$\mathcal{L}\simeq\mathcal{L}_{6}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=0.$
-$\mathcal{L}\simeq\mathcal{L}_{7}^{\lambda\neq 0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=(1+2\lambda)\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$
-$\mathcal{L}\simeq\mathcal{L}_{7}^{0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=3.$
-$\mathcal{L}\simeq\mathcal{L}_{9}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$
+$\mathcal{L}\simeq\mathcal{L}_{1}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=2.$
+$\mathcal{L}\simeq\mathcal{L}_{2}^{\lambda\neq 0}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=(\lambda-1)\mathop{\rm tr}(\chi_{1}\chi_{0}).$
+$\mathcal{L}\simeq\mathcal{L}_{3}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=-\mathop{\rm tr}(\chi_{1}\chi_{0}).$
+$\mathcal{L}\simeq\mathcal{L}_{6}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=0.$
+$\mathcal{L}\simeq\mathcal{L}_{7}^{\lambda\neq 0}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=(1+2\lambda)\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$
+$\mathcal{L}\simeq\mathcal{L}_{7}^{0}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=3.$
+$\mathcal{L}\simeq\mathcal{L}_{9}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$
+$\mathcal{L}\simeq\mathcal{L}_{10}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=4.$
+Acknowledgement
+We would like to express our sincere gratitude to the anonymous referee for the valuable and constructive comments, which have helped us improve the quality and clarity of our paper.